### v1

# Good Noise Makes Good Edits: A Training-Free Diffusion-Based Video Editing with Image and Text Prompts

Abstract
We propose ImEdit, the first zero-shot, training-free video editing method conditioned on both images and text. The proposed method introduces $\rho$-start sampling and dilated dual masking to construct well-structured noise maps for coherent and accurate edits. We further present zero image guidance, a controllable negative prompt strategy, for visual fidelity. Both quantitative and qualitative evaluations show that our method outperforms state-of-the-art methods across all metrics.
Figure 1: Comparison results of video editing. ImEdit effectively edits source videos using image and text prompts, achieving high fidelity and temporal consistency without any additional training.
1 Introduction
Diffusion-based video editing methods have shown significant achievements in recent years, enabling precise modifications of objects in videos based on prompts provided by users stablevideo ; videdit ; videoswap ; genvideo ; videop2p ; dreamix ; revideo ; fatezero ; videoanydoor ; tune_a_video ; controlvideo ; make_a_pro ; cutandpaste . Primarily, text-driven video editing models have gained considerable attention for their ability to modify video content while maintaining temporal consistency tokenflow ; videop2p ; fatezero ; tune_a_video . These methods have demonstrated impressive results across a wide range of applications, such as appearance editing stablevideo ; videdit ; magicedit ; videop2p , and style transfer structure ; tokenflow ; rerenderavideo ; controlvideo .
However, text-based editing has inherent limitations, particularly when modifying objects with intricate and unique details, where textual descriptions alone are often insufficient to precisely convey the desired attributes precisely (Fig. 2(a)). Moreover, crafting detailed textual prompts is cumbersome, requiring extensive trial-and-error to achieve the intended modifications. This challenge becomes even more pronounced when the editing method necessitates one-shot finetuning for each edit videop2p ; fatezero ; save . A more efficient and intuitive alternative is image-driven prompting, which provides direct visual guidance for more precise and straightforward modifications.
Despite its advantages, image-driven video editing, particularly training-free approaches, remains largely unexplored. Previous work, Make-A-Protagonist make_a_pro , introduces an approach that trains a text-to-video (T2V) diffusion model conditioned on both text and image inputs via CLIP encoders clip . Similarly, GenVideo genvideo proposes a mask generation approach using an image diffusion model to address shape and size mismatches between the source object in the source video and the target object in the reference image. While effective, these methods require one-shot fine-tuning for each video, making the editing process highly costly and sensitive to variations in input resolution.
To overcome these limitations, a training-free video editing approach that conditions on both image and text prompts is essential. However, a key challenge remains: maintaining temporal consistency while preserving motion dynamics is fundamental to video editing, but becomes particularly difficult in a training-free setting, leading to color shifts, shape distortions, or loss of fine details. As shown in Fig. 2(b), simply integrating an image prompt into an existing T2V model without fine-tuning results in significant quality degradation, particularly in object shape fidelity.
Figure 2: Importance of appropriate integration of image prompts in video editing. Given (a) a source video, we compare (b) a state-of-the-art text-driven video editing method, (c) a naive combination of a pretrained T2V model with an image-guided module, and (d) our proposed method. The text-only approach in (b) struggles to accurately transfer the fine-grained details of the car in the reference image. Simply attaching the image-guided module to a pretrained T2V model, as in (c), produces unnatural results. In contrast, (d) our proposed training-free method precisely transfers the car in the reference image into the video with high fidelity.
Therefore, we propose that well-structured noise maps, generated through a controlled combination of the source video’s latent representation and Gaussian noise, effectively preserve motion consistency even without additional training. These noise maps act as a key component in achieving high-quality video edits. To further refine the noise maps to ensure seamless object transitions, we introduce a dilated dual masking strategy that blends the target object with the original video, which becomes particularly effective in handling shape discrepancies between the source and target objects. We also observe that a simple yet effective enhancement to the existing negative image prompt formulation used in IP-Adapter ipadapter effectively leads to realistic edits without requiring additional training.
Building on these findings, we present ImEdit, the first zero-shot, training-free video editing approach conditioned on both image and text prompts, enabling high-fidelity object replacement while preserving temporal consistency. Our method leverages well-structured craft noise maps for motion preservation, dilated dual masking for seamless object replacement, and zero image guidance for color fidelity. Additionally, Our approach is highly flexible to varying image resolutions, by leveraging a T2V backbone pretrained on diverse resolutions without additional fine-tuning, ensuring robust generalization across different input scales.
As illustrated in Fig. 2(c), our proposed method effectively modifies video objects based on reference images and simple text prompts, eliminating the need for complex prompt engineering. We also evaluate the time efficiency of the proposed method against baseline methods in Fig. 3. Across all tested input video sizes, our method consistently achieves the fastest runtime, demonstrating its advantage as a training-free solution for video editing. Notably, Fig. 1 highlights that our method effectively addresses challenging scenarios, including small source objects (top), significant color differences between the source and target objects (bottom left), and dynamic motion in the source video (bottom right), while maintaining motion consistency throughout the edited results. In contrast, baseline methods often fail to accurately reflect the reference appearance under these settings.
Figure 3: Comparison of runtime across different methods. Results are averaged over three runs for each unique input video size, with standard deviation.
Through extensive quantitative and qualitative comparisons, we demonstrate that our approach outperforms existing state-of-the-art video editing methods in both visual quality and temporal stability. By eliminating the need for fine-tuning, our method is efficient, scalable, and widely applicable across various video editing tasks. In summary, our key contributions are as follows:
  • We propose ImEdit, the first zero-shot, training-free video editing approach that leverages both text and image prompts to seamlessly replace objects while preserving original motion dynamics and temporal consistency.
  • We introduce an effective noise map construction strategy to guide the denoising U-Net for coherent and accurate video edits. As key components, $\rho$-start sampling preserves motion patterns and spatial layout of the source video, while dilated dual masking ensures smooth transitions in appearance and shape.
  • To further enhance color saturation and visual fidelity, we present zero image guidance, a controllable negative image prompt strategy that blends a zero image embedding into the existing negative image guidance.
  • Comprehensive quantitative and qualitative evaluations demonstrate that our method outperforms existing state-of-the-art approaches.
Figure 4: Overview of our proposed training-free approach. ImEdit constructs two levels of noise maps to guide the editing process. First, a rough noise map is constructed from the source video and denoised using a pretrained T2V model with image and text guidance, producing a rough edit. Next, a refined good noise map is constructed and denoised under the same conditions to produce the final video. In both stage, we denoise for $\rho$ steps by employing $\rho$-start sampling. The output more faithfully reflects the input prompts while preserving motion and scene structure.
2 Methods
We introduce ImEdit, a training-free video editing approach that performs object replacement while preserving the overall coherence of the scene, guided by text or image prompts. Our method is grounded in the careful integration of well-structured noise with appropriately chosen components. Specifically, we consider two key ingredients: the foreground latent, obtained via $\rho$-start sampling, and the background latent, acquired through DDIM inversion of the source video. These latents are blended using a carefully designed mask derived from both the source and target object masks, employing dilated dual masking to ensure smooth transitions and spatial consistency. In addition, we apply zero image guidance, a negative prompt-based strategy, to enhance color saturation and overall visual fidelity in the final output.
2.1 Preliminaries
Latent diffusion models (LDMs)
In diffusion-based image generation models, LDMs have become a standard approach due to their computational efficiency and ability to generate high-resolution, photorealistic images dalle2 ; stablediffusion ; imagen . These models leverage a pretrained autoencoder vae , consisting of an encoder $\mathcal{E}(\cdot)$ and a decoder $\mathcal{D}(\cdot)$, to map input images into a lower-dimensional latent space. Given an input image ${\bm{x}}$, the clean latent variable is obtained as ${\bm{z}}_{0}=\mathcal{E}({\bm{x}})$. During the forward process, gaussian noise is gradually added to the latent variable over $T$ steps. In the reverse process, a denoising network $\epsilon(\cdot)$ predicts noise at each step to iteratively refine the latent variable. Once the denoising process is complete, the decoder $\mathcal{D}(\cdot)$ transforms the latent variable back into the image space to generate the final output.
DDIM sampling and inversion
[…]
DDIM inversion predicts the sequence of noisy latents $\{{\bm{z}}_{t}\}_{t=1}^{T}$ from a given clean sample ${\bm{z}}_{0}$, reconstructing the corresponding diffusion trajectory without requiring any model retraining ddim ; inverting . We denote this inversion trajectory as DDIM-inv$({\bm{z}}_{0})$.
2.2 Overview
Given a source video $\mathcal{V}_{s}\in\mathbb{R}^{F\times C_{s}\times H_{s}\times W_{s}}$, our goal is to edit the foreground object and generate an edited video $\mathcal{V}_{e}\in\mathbb{R}^{F\times C_{s}\times H_{s}\times W_{s}}$, guided by a reference image $\mathcal{R}\in\mathbb{R}^{C_{r}\times H_{r}\times W_{r}}$ and a text prompt $\mathcal{T}$, while preserving the original background and motion dynamics. Here, $F$ is the number of frames, $C_{s}$ is the number of source video channels, and $H_{s}$ and $W_{s}$ are the height and width of each frame. Similarly, $C_{r},H_{r},$ and $W_{r}$ represent the channel and spatial dimensions of the reference image.
Our approach achieves this goal through a two-stage editing pipeline, as illustrated in Fig. 4. In the first stage, we construct a rough noise to enable localized object modification. In the second stage, this rough noise is refined into a more well-structured good noise, leading to higher-fidelity editing.
2.3 How to construct a good noise for video editing?
We first encode the source video $\mathcal{V}_{s}$ into the latent space using a pretrained VAE encoder $\mathcal{E}$. We then initialize the starting noisy latent using $\rho$-start sampling (see Section 2.3.2), a modification of the standard DDIM sampling schedule, which enables better preservation of the source video’s structure and motion. This noisy latent is passed into a pretrained T2V diffusion model $\epsilon_{\theta}$ where we incorporate both image and text guidance using the IP-Adapter ipadapter . The image and text features from CLIP image encoder $\text{CLIP}_{\text{img}}$ and CLIP text encoders $\text{CLIP}_{\text{txt}}$ clip are injected into the each corresponding cross-attention layers of the T2V model as condition embeddings:
${\bm{c}}_{\text{img}}=\text{CLIP}_{\text{img}}(\mathcal{R})\quad\text{ and }\quad{\bm{c}}_{\text{txt}}=\text{CLIP}_{\text{txt}}(\mathcal{T}),$ (2)
In addition, negative guidance using $\bar{{\bm{c}}}_{\text{img}}$ and $\bar{{\bm{c}}}_{\text{txt}}$, which denote the negative image and text embeddings respectively, helps suppress irrelevant visual or textual features.
During the denoising process with condition embeddings, we employ a foreground-background separation mechanism to perform object-centric modifications while preserving background fidelity, inspired by latent blending latentblending ; blendeddiffusion ; videoswap ; genvideo . Specifically, we split the noisy latent ${\bm{z}}_{t}$ into foreground and background regions using the object segmentation mask ${\bm{M}}$. For the background region, we retrieve the DDIM inversion noise corresponding to the current time step to preserve the background of the source video $\mathcal{V}_{s}$. In contrast, the foreground region is denoised by iteratively updating the noisy latent from the previous time step using a denoising network $\epsilon_{\theta}$, guided by the condition embeddings. This enables the foreground object to be progressively edited to resemble the target object. Formally, the foreground latent ${\bm{z}}_{t}^{\text{{fg}}}$ and background latent ${\bm{z}}_{t}^{\text{{bg}}}$ at time step $t$ are computed as:
[…]
${\bm{z}}^{\text{{bg}}}_{t\texttt{-}1}$ $={\bm{z}}_{t\texttt{-}1},\ \text{with }{\bm{z}}_{t-1}\in\{{\bm{z}}_{t}\}_{t=1}^{\tau_{\nu}}\text{ from }\text{DDIM-inv}({\bm{z}}_{0}).$ (4)
We then blend these two latents using ${\bm{M}}$ via our dilated dual masking (Section 2.3.1), as follows:
${\bm{z}}^{\prime}_{t\texttt{-}1}={\bm{z}}^{\text{{fg}}}_{t\texttt{-}1}\odot{\bm{M}}+{\bm{z}}^{\text{{bg}}}_{t\texttt{-}1}\odot(1-{\bm{M}}),$ (5)
where $\odot$ denotes element-wise multiplication. We use different mask $M$ at each stage. Specifically, in the first stage, the dilated source mask from the source video, ${\bm{M}}^{\prime}_{s}\in\mathbb{R}^{F\times C_{s}\times H_{s}\times W_{s}}$, is used. Through the denoising process in the first stage, we obtain a roughly edited video $\mathcal{V}_{r}$, which reflects the general structure and appearance of the target object. However, this rough edit may still retain residual features of the source object, which should be removed for a clean transformation.
To address the limitations of single mask blending, we introduce a second stage that further refines the output. In this stage, we generate the final mask ${\bm{M}}_{e}$ by combining the initial mask ${\bm{M}}^{\prime}_{s}$ with the rough mask ${\bm{M}}^{\prime}_{r}$ derived from the rough edit. Foreground latent ${\bm{z}}_{t}^{\text{{fg}}}$ and background latent ${\bm{z}}_{t}^{\text{{bg}}}$ are regenerated from the rough edit $\mathcal{V}_{r}$ using the same procedure as in the first stage. With the final mask and the regenerated latents, we construct a more refined noise map, which we refer to as good noise. It is designed to erase the original object and accurately follow the shape of the new object while maintaining the background. After denoising steps in the second stage, latents are decoded using the pretrained VAE decoder $\mathcal{D}$ to generate the final output. Note that the entire pipeline is training-free, relying solely on pretrained models for mask extraction and video generation.
Figure 5: Effective noise construction using our dilated dual masking. We blend the foreground and background latents, obtained through our proposed methods such as $\rho$-start sampling and zero image guidance, using a final mask constructed by taking the union of the dilated source and target masks. This enables clean removal of the source object and seamless replacement with the target.
2.3.1 Dilated dual masking
The effectiveness of our method stems from the careful coordination of structured noise using the proposed dilated dual mask, which plays a crucial role in defining the editable regions within each frame. This mask is central to enabling selective editing of target objects, as illustrated in Fig. 5.
In the first stage, only the source object mask ${\bm{M}}_{s}$ is available, which roughly outlines the foreground region to be edited. However, this mask lacks contextual adaptation to the background, often resulting in sharp edges and unnatural transitions at the object boundaries. Thus, we apply morphological dilation to smooth boundaries and facilitate seamless object replacement. Specifically, a single-iteration morphological dilation is applied to a binary matrix $M$ with a kernel of size $k\times k$. For each pixel position $(p,q)$, the output value is determined as the maximum value within a neighborhood defined by ${\mathcal{N}}_{k}$, where ${\mathcal{N}}_{k}=\{(i,j)\mid-r\leq i,j\leq r\}$ and $r=\lfloor\frac{k}{2}\rfloor$. The dilation operation is formally expressed as follows:
${{\bm{M}}^{\prime}}[p,q]=\max_{(i,j)\in{\mathcal{N}}_{k}}{\bm{M}}[p+i,q+j].$ (6)
As a result, the dilated source mask ${\bm{M}}^{\prime}_{s}$ is used to blend latents in Eq. 5 to produce rough edits $\mathcal{V}_{r}$.
In the second stage, we similarly obtain the dilated rough mask ${\bm{M}}^{\prime}_{r}$ from the rough edits. However, blending foreground and background latents using either the source mask ${\bm{M}}^{\prime}_{s}$ or the rough mask ${\bm{M}}^{\prime}_{r}$ poses the risk that parts of the source video may remain visible or that the generation area for the target object may not be fully covered. To address this, we leverage masks from both the first and second stages. The final mask ${\bm{M}}_{e}$ is computed as the union of the source mask ${\bm{M}}^{\prime}_{s}$ and the rough mask from the first stage ${\bm{M}}^{\prime}_{r}$, as shown in the following equation:
${\bm{M}}_{e}={\bm{M}}^{\prime}_{s}\cup{\bm{M}}^{\prime}_{r}.$ (7)
Finally, we construct the good noise ${\bm{z}}^{\prime}_{t}$ using the final mask ${\bm{M}}_{e}$ to guide the final blending process in Eq. 5. This enables the second stage to generate more realistic, artifact-free results, as ${\bm{M}}_{e}$ precisely defines the edited region. In particular, it ensures clean removal of the original object, accurate reconstruction of the new object’s shape, and consistent preservation of the background.
2.3.2 $\mathbf{\rho}$-start sampling
Preserving consistency with the source video is critical in video editing, particularly for maintaining visual structure and motion save ; camel ; deco . We hypothesize that initiating the sampling process from a less perturbed latent, closer to the source video, facilitates the preservation of these characteristics.
To this end, we propose an alternative sampling strategy that modifies the standard DDIM schedule by starting the denoising process from a less noisy latent representation, thereby improving structural and temporal fidelity. In the conventional reduced DDIM schedule ddim , sampling proceeds along a trajectory ${{\bm{z}}_{\tau_{1}},\ldots,{\bm{z}}_{\tau_{\nu}}}$, starting from a highly noisy latent ${\bm{z}}_{\tau_{\nu}}$. In contrast, our approach begins from an earlier timestep, ${\bm{z}}_{\tau_{\rho}}$, where $\rho<\nu$. This shortens the denoising trajectory to only $\rho$ steps, $\{{\bm{z}}_{\tau_{1}},\ldots,{\bm{z}}_{\tau_{\rho}}\}$, facilitating the retention of more information from the source video.
The parameter $\rho$ controls how early the sampling begins in the reverse diffusion process. A smaller $\rho$ initiates from a latent closer to the source video, enabling better preservation of motion patterns and spatial layout, such as object direction and position. Conversely, a larger $\rho$ facilitates greater transformation, allowing the edited video to more closely resemble the reference image in terms of object appearance. By adjusting $\rho$, we gain fine-grained control over the degree of signal retention, enabling higher fidelity in both motion and overall scene consistency.
2.4 Zero image guidance
To enhance color fidelity, we introduce zero image guidance, helping better align with the appearance of the reference image. Our refined guidance strategy blends the conventional negative image embeddings CFG ; glide , represented by a zero vector of the same size as the CLIP image embedding, with a zero image guidance, which refers to the CLIP image embedding derived from an all-zero image. The final negative image guidance $\bar{{\bm{c}}}_{\text{img}}$ is computed as a weighted sum of the two, controlled by a blending factor $\gamma$, and can be formally expressed as:
$\bar{{\bm{c}}}_{\text{img}}=\gamma\cdot E_{I}({\mathbf{I}}_{0})+(1-\gamma)\cdot\mathbf{0},$ (8)
where $\gamma\in[0,1]$ is a hyperparameter, $\mathbf{0}\in\mathbb{R}^{d}$ is a zero vector with the same size as the CLIP image embedding, $E_{I}(\cdot)$ represents the CLIP image encoder, and ${\mathbf{I}}_{0}\in\mathbb{R}^{C\times H\times W}$ is an all-zero image input. This formulation enables a tunable trade-off between the two types of negative guidance, allowing finer control over the color intensity of the generated object. The resulting $\bar{{\bm{c}}}_{\text{img}}$, together with the conventional $\bar{{\bm{c}}}_{\text{txt}}$, serves as negative guidance during the sampling process in Eq. 3.
3 Experiments
3.1 Experimental setup
Evaluation metrics. We conduct evaluations on 18 pairs of video and reference image, using videos collected from DAVIS dataset davis and the dataset released by Video-P2P videop2p . To assess how well the generated videos align with the given text and image prompt, we use CLIP-T score and DINO score. Specifically, CLIP-T score measures the alignment between the target text prompt and the generated videos using CLIP text encoder clip , while DINO score calculates the similarity between the reference image and the generated video using DINO dino . To evaluate temporal consistency, we extract CLIP image embeddings for each frame in the edited video using the CLIP image encoder clip and measure the cosine similarity between embeddings of adjacent frames.
Baseline models. We compare our proposed method with two baselines: Video-P2P videop2p and Make-A-Protagonist (Make-A-Pro) make_a_pro . To the best of our knowledge, Make-A-Pro is the only baseline that directly aligns with our problem setting and offers publicly released code among existing methods. However, in order to comprehensively demonstrate the effectiveness of our image prompt approach, we also include Video-P2P, a state-of-the-art text-driven video editing model, as an additional baseline.
Implementation details. We use AnimateDiff animatediff as the base T2V model, IP-Adapter ipadapter v1.5 for the reference image encoding, and SAM2 sam2 for obtaining masks. We conduct experiments with videos of various resolutions (320$\times$576, 480$\times$480, 512$\times$512, 600$\times$600 and 640$\times$640) and different numbers of frames (8 and 16). For text prompts, we use the same descriptions when comparing our method with Video-P2P videop2p . In contrast, for Make-A-Pro make_a_pro , we follow their strategy by extracting protagonist information using BLIP-2 blip2 and generating prompts via visual question answering. Additionally, since Video-P2P videop2p only supports square outputs, we paste its generated video back into the corresponding regions of the source video to enable a comparison. All experiments are conducted on a single NVIDIA A100 GPU with 40GB memory.
3.2 Experimental results
3.2.1 Comparisons with state-of-the-art methods
Quantitative and qualitative results. We quantitatively evaluate our method with a focus on alignment with text and image prompts, visual feature similarity, and temporal consistency across video frames. As shown in Table 1, our method consistently outperforms all baselines across all evaluation metrics, both automatic and human. In terms of model scoring metrics, our method achieves the highest CLIP-T, DINO, and Temp scores, indicating superior text alignment, object similarity, and temporal consistency, respectively. These trends are further supported by human evaluation. Our method receives the lowest scores in Image, Text, and Temp categories, demonstrating that it generates objects most similar to the target, aligns most faithfully with the input text prompts, and maintains the most consistent appearance across frames. In addition, our method is rated highest in Overall quality, clearly confirming its effectiveness and perceptual superiority over the baselines.
Qualitative comparisons further demonstrate the effectiveness of our method in accurately preserving both motion and identity across diverse editing scenarios, outperforming the baselines. In the left example of Fig. 6, Make-A-Pro make_a_pro preserves the motion from the source video but fails to accurately reflect the identity of the reference image, even struggling to edit the dog into a fox. Video-P2P videop2p partially captures the identity of the fox but fails to preserve motion, resulting in an incorrect number of objects, with two foxes appearing in a single frame. In contrast, our method produces the most accurate edited video, successfully maintaining both the motion from the source video and the identity of the reference object. The right example in Fig. 6 features a source video with less dynamic movement. In this case, all methods preserve motion, but only our method correctly generates a white duck as specified in both the text condition and reference image. This example highlights the ability of our method to effectively handle both image and text prompts in video editing tasks.
Table 1: Quantitative comparison results. Quantitative metrics include CLIP-Text similarity (CLIP-T), visual feature similarity (DINO), and temporal consistency (Temp). Human evaluations measure the text-video alignment (Text), visual fidelity to the target object (Image), motion consistency with the source video (Motion), temporal consistency (Temp), and overall visual quality (Overall).
Method Model scoring metrics Human evaluations
CLIP-T $\uparrow$ DINO $\uparrow$ Temp $\uparrow$ Text $\downarrow$ Image $\downarrow$ Motion $\downarrow$ Temp $\downarrow$ Overall $\downarrow$
Video-p2p videop2p 0.300 0.393 0.961 2.01 2.05 2.27 1.99 2.17
Make-A-Pro make_a_pro 0.294 0.362 0.958 2.66 2.61 1.99 2.64 2.57
ImEdit (Ours) 0.326 0.474 0.966 1.33 1.34 1.74 1.37 1.26
Figure 6: Qualitative comparison results on DAVIS dataset. We compare the generation quality of our method against baseline methods. While the baseline methods fail to incorporate key visual features from the reference image, our method effectively transfers the target object into the source video, faithfully preserving its appearance across frames.
User study. We conduct a user study based on five evaluation criteria: (1) image alignment, which measures how well the object in the generated video resembles the one in the reference image; (2) text alignment, which assesses how accurately the generated video matches the given text condition; (3) motion alignment, which evaluates how well the motion in the generated video aligns with that of the source video; (4) temporal consistency, which examines the consistency of the object and background across frames; and (5) overall quality, which considers all the aforementioned factors to determine how natural and realistic the generated video appears. We ask 32 participants to evaluate 9 comparisons each, ranking the given results based on these criteria. We compute the average rankings, and Table 1 shows that our method ranks highest across all criteria.
3.2.2 Ablation study
Quantitative results. We quantitatively evaluate the contribution of each component in our method, as shown in Table 2. Starting from the base model (a), the integration of zero image guidance (b) leads to an overall performance improvement. Incorporating $\rho$-start sampling (c) brings a slight increase in CLIP-Text alignment, while the DINO and Temporal scores slightly decrease, indicating that adjusting the starting sampling point alone is insufficient to resolve the trade-off between motion preservation and appearance fidelity. The inclusion of dual masking (d) contributes to a more substantial gain across all metrics. Finally, our proposed method with dilation (e) achieves the best performance, with notable improvements. These results highlight the effectiveness of the complete method.
Effectiveness of dilated dual masking. We ablate our proposed dilated dual masking by progressively incorporating each masking schemes, as shown in Fig. 7. We begin with the simplest configuration: a single stage editing setup that relies solely on the source mask. When using the source mask only, the generated object is often forced to conform to the contours of the source object, leading to unrealistic shapes, limited fidelity, and artifacts. This issue becomes particularly pronounced when there is a significant shape difference between the source and target objects. As shown in Fig. 7(a), the source object (grey car) is structurally different from the target object; it is shorter in length, taller in height, and more angular than the target object, but these differences are not considered.
Adding an additional stage to generate the rough mask and using only this mask alleviates the problem of the imprecise reflection of the target object’s features. However, Fig. 7(b) reveals that it fails to completely remove the source, as the top part of the grey car still remains. To further address this, we take the union of the source and rough mask. While this reduces residuals from the source object, artifacts persist along the boundaries. Furthermore, fine-grained details of the generated target object are lost due to the strict constraints imposed by the merged mask region, as illustrated in Fig. 7(c).
Our proposed dilated dual masking effectively addresses the aforementioned issues by applying dilation to the masks and this is evident in Fig. 7(d). It allows the generated target object to extend beyond the strict union region, enabling natural shape adaptation and smoother blending into the surrounding context. As a result, the proposed dilated dual masking achieves both structural integrity and visual realism, demonstrating its superiority.
Table 2: Comprehensive ablation study of ImEdit. We evaluate the contribution of each component to the overall performance of our method. (a) Base model. (b) Zero image guidance. (c) $\rho$-start sampling. (d) Dual masking. (e) Dilation.
Method CLIP-Text $\uparrow$ DINO $\uparrow$ Temp $\uparrow$
(a) 0.302 0.361 0.958
(a) + (b) 0.304 0.405 0.960
(a) + (b) + (c) 0.309 0.401 0.955
(a) + (b) + (c) + (d) 0.316 0.435 0.958
(a) + (b) + (c) + (d) + (e) (Ours) 0.326 0.474 0.966
Figure 7: Ablation study on the proposed dilated dual masking. (a) shows the output of the first stage using the source video $\mathcal{V}_{s}$ and source object mask, while (b)-(d) present outputs of the second stage from the intermediate video $\mathcal{V}_{r}$ using different masking strategies. Masks are overlaid on the corresponding video. The proposed dilated dual masking (d) achieves the most accurate edits, underscoring its effectiveness for the object replacement.
Figure 8: Effect of varying the $\rho$ value. The rightmost column shows the final output after the second stage, while all other columns, excluding the reference image and source video, show the results of the first stage under different values of $\rho$. By adjusting sampling start point, we balance how strongly the appearance of the target object is retained versus the source motion. Lower $\rho$ emphasizes the owl’s visual traits, while higher $\rho$  increasingly preserves the parrot’s original movement and posture.
Effectiveness of $\rho$-start sampling. As shown in Fig. 8, varying the parameter $\rho$, which determines the sampling start point, enables control over the balance between the target appearance and source motion. Using a larger $\rho$ leads the model to fails to accurately follow the motion and spatial layout of the source video. For example, while the parrot in the source video is facing left, the owl in the edited video is facing forward or at an incorrect angle. In contrast, using a smaller $\rho$ allows the model to better capture the motion and spatial dynamics of the source video, resulting in the owl correctly facing sideways, consistent with the source object. However, if $\rho$ is too small, the transformation may become too weak, leading to incomplete replacement of the source object. With an appropriate choice of $\rho$  the edited video successfully balances appearance fidelity and motion preservation, highlighting the importance of temporal alignment in video editing tasks.
Effectiveness of the zero image guidance. We analyze the impact of $\gamma$ in the proposed zero image guidance. As shown in Fig. 9, increasing $\gamma$, which scales the contribution of the zero image guidance (see Eq. (8)), leads to more saturated colors. This suggests that mixing the zero image guidance into the negative image prompt effectively enhances color intensity, providing a simple yet powerful control for visual intensity.
Figure 9: Effect of varying the $\gamma$ value. Increasing $\gamma$ enhances the influence of the zero image embedding, leading to more saturated color reproduction in the edited video.
4 Conclusion
In this work, we propose ImEdit, a novel zero-shot, training-free video editing method guided by both image and text prompts. ImEdit enables accurate object replacement with strong temporal coherence and achieves superior performance in both visual quality and motion consistency, outperforming existing state-of-the-art approaches with high efficiency and scalability.
Appendix
This supplementary material complements the main manuscript, providing additional experimental results and implementation details.
  • Appendix A provides an analysis of the effect of the DDIM inversion trajectory length.
  • Appendix B provides an additional quantitative comparison with baseline methods.
  • Appendix C provides related work.
  • Appendix D addresses the limitations of our method.
  • Appendix E provides additional ablation results on dilated dual masking.
  • Appendix F provides additional implementation details and user study details.
  • Appendix G presents the overall algorithm of the proposed method.
  • Appendix H provides additional qualitative results.
Appendix A Effect of the DDIM inversion trajectory length
In our latent blending process, the number of DDIM inversion sampling steps is treated as a hyperparameter with two discrete settings, 999 or 50, depending on the desired outcome. Using 999 steps often results in less vivid colors and a relatively noisy background. However, the edited object tends to exhibit more natural shapes and the overall video appears more realistic. In contrast, using 50 steps yields cleaner, more vivid appearances for both the object and background, and significantly reduces DDIM inversion time. Nonetheless, the shape of the edited object may appear less natural.
As shown in Fig. 10, the result with 999 steps shows a polar bear with reduced color saturation and increased background noise while the head and front paws are generated with a more natural structure. Conversely, the result with 50 steps shows a polar bear with more vivid white appearance and cleaner background, but the shape of the head and paws is less natural.
Figure 10: Effect of the reduced inversion trajectory length. We evaluate how varying $\nu$, the number of DDIM steps used in inversion, affects editing quality.
Appendix B Additional quantitative results
Table 3 presents an additional quantitative comparison between ImEdit and existing baseline methods using CLIP-T, DINO, Temp, and PSNR, which respectively assess text-image alignment, the similarity between the reference image and the edited video, temporal consistency, and background preservation. To facilitate a more comprehensive evaluation, we extend the comparison in Table 1 by including FateZero fatezero and a naive variant of our approach, AnimateDiff animatediff combined with IP-Adapter ipadapter , and by adding PSNR as an additional metric. Specifically, AnimateDiff + IP-Adapter refers to the direct combining of our backbone model AnimateDiff with IP-Adapter, without any of the proposed components, as illustrated in Fig. 2. FateZero fatezero is included as a baseline due to its strong editing performance among text-driven video editing methods that operate in a training-free setting, consistent with the setting of our method.
Table 3: Additional quantitative comparison with baseline models. We compare our method (ImEdit) against existing baselines across four metrics: CLIP-Text similarity (CLIP-T), visual feature similarity (DINO), temporal consistency (Temp), and background preservation (PSNR).
Method Model scoring metrics
CLIP-T $\uparrow$ DINO $\uparrow$ Temp $\uparrow$ PSNR $\uparrow$
FateZero fatezero 0.317 0.460 0.960 21.32
Video-p2p videop2p 0.300 0.393 0.961 21.75
Make-a-Pro make_a_pro 0.294 0.362 0.958 15.40
AnimateDiff animatediff + IP-Adapter ipadapter 0.302 0.361 0.958 24.83
ImEdit (Ours) 0.326 0.474 0.966 23.59
As shown in Table 3, ImEdit outperforms Video-P2P videop2p , Make-a-Pro make_a_pro , and FateZero fatezero across all metrics. Notably, in terms of DINO scores, our approach achieves higher reference similarity not only compared to text-based methods such as Video-P2 videop2p and FateZero fatezero but also surpasses image-conditioned methods like Make-a-Pro make_a_pro . While AnimateDiff + IP achieves a higher PSNR than ours, this difference stems from the smaller edited region in its outputs. ImEdit employs dilated dual masking during editing, which results in a broader generation area and a slightly lower PSNR.
Appendix C Related work
C.1 Image and Video Generation
Recent advancements in latent diffusion models (LDMs) have significantly improved the fidelity and controllability of image generation. Text-to-image (T2I) models such as Stable Diffusion stablediffusion , DALL-E 2 dalle2 , and Imagen imagen achieve high-resolution image synthesis aligned with complex text prompts. These models operate in the latent space for computational efficiency and support diverse conditioning mechanisms, including text, layout, and edge maps ipadapter ; controlnet . Their flexibility has made LDMs foundational components for extending generative modeling into the video domain.
Motivated by T2I success, text-to-video (T2V) models have emerged to generate coherent video sequences conditioned on textual descriptions. CogVideo cogvideo adapts T2I backbones with temporal attention, enabling efficient video generation with minimal overhead. ZeroScope zeroscope , built on ModelScope modelscope , improves visual fidelity and consistency, producing higher-resolution outputs while mitigating common artifacts in earlier methods. AnimateDiff animatediff integrates temporal modules into Stable Diffusion stablediffusion , facilitating the T2V model with improved coherence between frames.
Alongside architectural advances, controllability and identity preservation have gained importance. Several methods leverage a reference frame along with a prompt to guide spatial appearance and layout across time stiv ; ti2v ; controlvideo . In addition, others introduce subject-driven personalization via instance finetuning videoalchemist ; dreamvideo ; tune_a_video ; customcrafter , enabling video generation with identity consistency across diverse prompts. These trends illustrate a broader shift toward multimodal conditioning across text, image, and motion to allow more expressive and controllable video synthesis.
C.2 Image and Video Editing
Following the advances in image and video generation models, diffusion models have also made significant progress in image and video editing, allowing intuitive and precise content manipulation. In the image domain, Prompt-to-Prompt prompt-to-prompt introduces cross-attention manipulation within pretrained diffusion models, enabling semantic edits without finetuning or masks. InstructPix2Pix instructpix2pix leverages pairs of instructions and images to support editing guided by natural language. EditAnything editanything generalizes this paradigm by integrating multi-modal inputs such as text, masks, and user clicks.
Video editing extends these capabilities to the temporal dimension. Tune-A-Video tune_a_video employs one-shot tuning of pretrained image diffusion models for video editing tasks, allowing users to modify video content and style with minimal data and strong temporal consistency. Video-P2P videop2p extends cross-attention control techniques from image to video, maintaining both spatial and temporal coherence. Text2Video-Zero text2video removes the need for paired text-video data, instead leveraging pretrained text-image models for flexible and data-efficient video editing.
To overcome the limitations of editing based solely on text, recent studies propose image-conditioned video editing methods that use a reference image to guide appearance and identity genvideo ; make_a_pro ; cutandpaste . Make-A-Protagonist make_a_pro introduces a unified framework for protagonist replacement, background editing, and style transfer using mask-guided denoising. GenVideo genvideo proposes shape-aware masking and latent noise correction to ensure temporal consistency and subject fidelity. VideoSwap videoswap adds semantic point correspondences to provide interactive and precise control over the subject’s motion and appearance. Despite these advancements, existing methods still require a finetuning stage with the source video, which limits their overall effectiveness. In contrast, our approach eliminates this need entirely, enabling high-quality, zero-shot video editing with strong temporal and identity coherence.
Appendix D Limitation
Although our proposed method exhibits notable improvements over existing baselines, as evidenced by extensive experiments, it still faces certain limitations. When there is a substantial disparity in the innate physical characteristics between the source and target objects, our method struggles to perform object replacement. Figure 11 shows that our method generates wings for the red panda which is an anatomically implausible feature to transfer the parrot’s flying motion. Conversely, in attempting to preserve the red panda’s natural characteristics, the method fails to convey the intended motion.
Figure 11: Example of failure case. A large semantic gap between the object in the source video and the reference image poses challenges for our method. We show results from two versions of our method, obtained by varying the sampling start point $\rho$ and the dilation value. In this case, editing a bird into a red panda results in missing or incorrect transfer of the wings and arms.
Figure 12: Additional results on the proposed dilated dual masking. (a) shows the output of the first stage using the source video $\mathcal{V}_{s}$ and source object mask, while (b)-(d) present outputs of the second stage from the intermediate video $\mathcal{V}_{r}$ using different masking strategies. Masks are overlaid on the corresponding video. The proposed dilated dual masking (d) achieves the most accurate edits, underscoring its effectiveness for the object replacement.
Appendix E Additional ablation results on dilated dual masking
The proposed dilated dual masking enables the seamless replacement of the source object with the target object from the reference image across all frames. Fig. 12 presents the complete results of Fig. 7, confirming that the patterns observed remain consistent across different frames.
In Fig. 12(a), when only the source mask is used, the target object is unnaturally constrained to the shape of the source mask, resulting in distorted generations. In Fig. 12(b), using only the rough mask fails to fully cover the source object, leaving undesired remnants of it in the edited video. In Fig. 12(c), employing a dual mask obtained by taking the union of the source and target masks alleviates artifacts from the source object and shape mismatches of the target object, but the mask boundaries still lack smoothness. In contrast, our proposed dilated dual mask in Fig. 12(d) effectively mitigates these issues, enabling natural removal of the source object and coherent synthesis of the target object throughout the video.
Appendix F Additional experimental settings
Appendix F.1 provides additional implementation details, including hyperparameter settings. Appendix F.2 describes the evaluation questions used in our user study.
F.1 Additional implementation details
Our method involves four key hyperparameters that control different aspects of the video editing process, which include $\rho,\gamma,\text{dilation kernel size $k$, and guidance scale}$:
  • $\rho\in[10,30]$: Control the starting timestep for the sampling process in $\rho$-start sampling, determining the trade-off between motion preservation from the source video and appearance transformation toward the reference image.
  • $\gamma\in[0,1]$: Control the weighted combination of standard zero-vector and zero-image embeddings in zero image guidance, enabling fine-grained adjustment of color fidelity in the generated object.
  • dilation kernel size $k$ $\in[0,10]$: Specifies the size of the dilation kernel used in the proposed dilated dual masking. This affects the spatial smoothness and coverage of the edited region.
  • guidance scale $\in[6,10]$: Determine the strength of classifier-free guidance during denoising. A higher guidance scale enhances fidelity to the conditioning signal.
F.2 User study details
In the user study described in Section  3.2.1, participants are asked to evaluate five key aspects of the generated videos using the following ranking-based questions:
  • Image Alignment: Rank the videos from 1st to 4th based on how well the appearance of the "object in the generated video" matches the "object in the reference image (excluding the background)".
  • Text Alignment: Rank the videos from 1st to 4th based on how well the "generated video content" aligns with the "provided text prompt".
    Text prompt: “a polar bear skiing”
  • Motion Alignment: Rank the videos from 1st to 4th based on how closely the "motion of the object in the generated video" matches the "motion of the object in the source video".
  • Temporal Consistency: Rank the videos from 1st to 4th based on the "visual consistency of the object and background throughout the edited video".
  • Overall Quality: Taking all the above aspects into consideration, rank the videos from 1st to 4th based on the overall realism and naturalness.
Appendix G Algorithm
The overall video editing procedure of ImEdit is summarized in Algorithm 1. Our training-free method consists of two stages: a rough editing stage that generates an initial object transformation guided by the source mask, and a refinement stage that further improves visual fidelity and temporal coherence using an updated mask derived from the intermediate result.
Algorithm 1 Training-free video editing procedure of ImEdit
1:Input: Source video $\mathcal{V}_{s}$, image prompt $\mathcal{R}$, text prompt $\mathcal{T}$, negative text prompt $\bar{\mathcal{T}}$
2:Output: Edited video $\mathcal{V}_{e}$
3: Prompt feature extraction
4:${\bm{c}}_{\text{img}}\leftarrow\text{CLIP}_{\text{img}}(\mathcal{R}),{\bm{c}}_{\text{txt}}\leftarrow\text{CLIP}_{\text{txt}}(\mathcal{T})$
5:$\bar{{\bm{c}}}_{\text{img}}=\gamma\cdot\text{CLIP}_{\text{img}}({\mathbf{I}}_{0})+(1-\gamma)\cdot\mathbf{0}$, $\bar{{\bm{c}}}_{\text{txt}}=\gamma\cdot\text{CLIP}_{\text{txt}}({\mathcal{\bar{T}}}),\text{ where }\gamma\in[0,1]$
6: Stage 1: Rough editing
7: Dilated mask generation
8:${\bm{M}}_{s}\leftarrow\text{Segment}(\mathcal{V}_{s})$
9:${\bm{M}}^{\prime}_{s}\leftarrow\text{Dilate}({\bm{M}}_{s})$
10: Latent encoding
11:${\bm{z}}_{0}\leftarrow\mathcal{E}(\mathcal{V}_{s})$
12:$\{{\bm{z}}_{t}\}_{t=\tau_{1}}^{\tau_{\nu}}\leftarrow\text{DDIM-inv}({\bm{z}}_{0})$
13: Denoising
14:Select $\rho\in[\nu]$
15:${\bm{z}}^{\prime}_{\tau_{\rho}}\leftarrow\sqrt{\alpha_{\tau_{\rho}}}\cdot{\bm{z}}_{0}+\sqrt{1-\alpha_{\tau_{\rho}}}\cdot\eta$, where $\eta\sim\mathcal{N}(0,{\bm{I}})$
16:for $t=\rho$ to $1$ do
17:     ${\bm{z}}^{\text{{fg}}}_{\tau_{t\texttt{-}1}}\leftarrow\text{DDIM-samp}({\bm{z}}^{\prime}_{\tau_{t}},{\bm{c}}_{\text{img}},{\bm{c}}_{\text{txt}},\bar{{\bm{c}}}_{\text{img}},\bar{{\bm{c}}}_{\text{txt}},\tau_{t})$
18:     ${\bm{z}}^{\text{{bg}}}_{\tau_{t\texttt{-}1}}\leftarrow{\bm{z}}_{\tau_{t\texttt{-}1}}$
19:     ${\bm{z}}^{\prime}_{\tau_{t\texttt{-}1}}\leftarrow{\bm{z}}^{\text{{fg}}}_{\tau_{t\texttt{-}1}}\odot{\bm{M}}^{\prime}_{s}+{\bm{z}}^{\text{{bg}}}_{\tau_{t\texttt{-}1}}\odot(1-{\bm{M}}^{\prime}_{s})$
20:end for
21: Latent decoding
22:$\mathcal{V}_{r}\leftarrow\mathcal{D}({\bm{z}}^{\prime}_{\tau_{0}})$
23:$\mathcal{V}_{r}\leftarrow\mathcal{V}_{r}\odot{\bm{M}}^{\prime}_{s}+\mathcal{V}_{s}\odot(1-{\bm{M}}^{\prime}_{s})$
24: Stage 2: Refined editing
25: Dilated dual mask generation
26:${\bm{M}}_{r}\leftarrow\text{Segment}(\mathcal{V}_{r})$
27:${\bm{M}}^{\prime}_{r}\leftarrow\text{Dilate}({\bm{M}}_{r})$
28:${\bm{M}}_{e}\leftarrow{\bm{M}}^{\prime}_{s}\cup{\bm{M}}^{\prime}_{r}$
29: Latent encoding
30:${\bm{z}}_{0}\leftarrow\mathcal{E}(\mathcal{V}_{r})$
31:$\{{\bm{z}}_{t}\}_{t=\tau_{1}}^{\tau_{\nu}}\leftarrow\text{DDIM-inv}({\bm{z}}_{0})$
32: Denoising
33:Select $\rho\in[\nu]$
34:${\bm{z}}^{\prime}_{\tau_{\rho}}\leftarrow\sqrt{\alpha_{\tau_{\rho}}}\cdot{\bm{z}}_{0}+\sqrt{1-\alpha_{\tau_{\rho}}}\cdot\eta$, where $\eta\sim\mathcal{N}(0,{\bm{I}})$
35:for $t=\rho$ to $1$ do
36:     ${\bm{z}}^{\text{{fg}}}_{\tau_{t\texttt{-}1}}\leftarrow\text{DDIM-samp}({\bm{z}}^{\prime}_{\tau_{t}},{\bm{c}}_{\text{img}},{\bm{c}}_{\text{txt}},\bar{{\bm{c}}}_{\text{img}},\bar{{\bm{c}}}_{\text{txt}},\tau_{t})$
37:     ${\bm{z}}^{\text{{bg}}}_{\tau_{t\texttt{-}1}}\leftarrow{\bm{z}}_{\tau_{t\texttt{-}1}}$
38:     ${\bm{z}}^{\prime}_{\tau_{t\texttt{-}1}}\leftarrow{\bm{z}}^{\text{{fg}}}_{\tau_{t\texttt{-}1}}\odot{\bm{M}}_{e}+{\bm{z}}^{\text{{bg}}}_{\tau_{t\texttt{-}1}}\odot(1-{\bm{M}}_{e})$
39:end for
40: Latent decoding
41:$\mathcal{V}_{e}\leftarrow\mathcal{D}({\bm{z}}^{\prime}_{\tau_{0}})$
Appendix H Additional qualitative results
Notably, AnimateDiff+IP-Adapter frequently exhibits noticeable distortions in both shape and color, leading to unnatural visual artifacts. To address this, our method constructs a well-formed noise maps by selecting a less-degraded latent via $\rho$-start sampling and combining foreground and background latents through dilated dual masking. This structured noise formulation enables accurate object replacement while preserving the motion dynamics of the source video.
Figure 13: Complete video editing comparisons with baseline methods. FateZero fatezero and Make-a-Pro make_a_pro use text prompts generated by BLIP-2 blip2 , following their respective proposed strategies. Video-P2P videop2p and FateZero fatezero use only text prompts, while Make-a-Pro make_a_pro , AnimateDiff animatediff +IP-Adapter ipadapter , and our proposed method use both image and text prompts.
Figure 14: Complete video editing comparisons with baseline methods. FateZero fatezero and Make-a-Pro make_a_pro use text prompts generated by BLIP-2 blip2 , following their respective proposed strategies. Video-P2P videop2p and FateZero fatezero use only text prompts, while Make-a-Pro make_a_pro , AnimateDiff animatediff +IP-Adapter ipadapter , and our proposed method use both image and text prompts.
Figure 15: Complete video editing comparisons with baseline methods. FateZero fatezero and Make-a-Pro make_a_pro use text prompts generated by BLIP-2 blip2 following their respective proposed strategies. Video-P2P videop2p and FateZero fatezero use only text prompts, while Make-a-Pro make_a_pro , AnimateDiff animatediff +IP-Adapter ipadapter , and our proposed method use both image and text prompts.

### v2

@@
 Abstract
-We propose ImEdit, the first zero-shot, training-free video editing method conditioned on both images and text.
-The proposed method introduces $\rho$-start sampling and dilated dual masking to construct well-structured noise maps for coherent and accurate edits.
-We further present zero image guidance, a controllable negative prompt strategy, for visual fidelity.
-Both quantitative and qualitative evaluations show that our method outperforms state-of-the-art methods across all metrics.
-Figure 1: Comparison results of video editing. ImEdit effectively edits source videos using image and text prompts, achieving high fidelity and temporal consistency without any additional training.
+We propose VINO, the first zero-shot, training-free video editing method conditioned on both image and text.
+Our approach introduces $\rho$-start sampling and dilated dual masking to construct structured noise maps that enable coherent and accurate edits.
+To further enhance visual fidelity, we present zero image guidance, a controllable negative prompt strategy.
+Extensive experiments demonstrate that VINO faithfully incorporates the reference image into video edits, achieving strong performance compared to state-of-the-art baselines, all without any test-time or instance-specific training.
 1 Introduction
-Diffusion-based video editing methods have shown significant achievements in recent years, enabling precise modifications of objects in videos based on prompts provided by users stablevideo ; videdit ; videoswap ; genvideo ; videop2p ; dreamix ; revideo ; fatezero ; videoanydoor ; tune_a_video ; controlvideo ; make_a_pro ; cutandpaste .
-Primarily, text-driven video editing models have gained considerable attention for their ability to modify video content while maintaining temporal consistency tokenflow ; videop2p ; fatezero ; tune_a_video .
-These methods have demonstrated impressive results across a wide range of applications, such as appearance editing stablevideo ; videdit ; magicedit ; videop2p , and style transfer structure ; tokenflow ; rerenderavideo ; controlvideo .
-However, text-based editing has inherent limitations, particularly when modifying objects with intricate and unique details, where textual descriptions alone are often insufficient to precisely convey the desired attributes precisely (Fig.
+Diffusion-based video editing methods have shown notable achievements in recent years, enabling precise modifications of objects in videos based on user prompts [stablevideo, videdit, dreamix, revideo, controlvideo, cutandpaste].
+Primarily, text-driven video editing models have gained considerable attention for their ability to modify video content while maintaining temporal consistency [tokenflow, videop2p, fatezero, tune_a_video].
+These methods have demonstrated impressive results across a wide range of applications, such as appearance editing [stablevideo, videdit, magicedit, videop2p], and style transfer [structure, tokenflow, rerenderavideo, controlvideo].
+Figure 2: Importance of appropriate integration of image prompts in video editing.
+Given (a) a source video, we compare (b) a state-of-the-art text-driven video editing method, (c) a naive combination of a pretrained T2V model with an image-guided module, and (d) our proposed method.
+The text-only approach in (b) struggles to accurately transfer the structural appearance of the car in the reference image.
+Simply attaching the image-guided module to a pretrained T2V model, as in (c), produces unnatural results.
+In contrast, our proposed training-free method in (d) precisely transfers the car in the reference image into the video with high fidelity.
+Figure 3: Runtime comparison across video resolutions. We compare VINO with Make-A-Pro [make_a_pro], an image-guided video editing method. Results are averaged over three runs per setting.
+However, text-based video editing has inherent limitations, particularly when modifying objects with intricate and unique details, where textual descriptions alone are often insufficient to precisely convey the desired attributes (Fig.
 2(a)).
 Moreover, crafting detailed textual prompts is cumbersome, requiring extensive trial-and-error to achieve the intended modifications.
-This challenge becomes even more pronounced when the editing method necessitates one-shot finetuning for each edit videop2p ; fatezero ; save .
-A more efficient and intuitive alternative is image-driven prompting, which provides direct visual guidance for more precise and straightforward modifications.
-Despite its advantages, image-driven video editing, particularly training-free approaches, remains largely unexplored.
-Previous work, Make-A-Protagonist make_a_pro , introduces an approach that trains a text-to-video (T2V) diffusion model conditioned on both text and image inputs via CLIP encoders clip .
-Similarly, GenVideo genvideo proposes a mask generation approach using an image diffusion model to address shape and size mismatches between the source object in the source video and the target object in the reference image.
-While effective, these methods require one-shot fine-tuning for each video, making the editing process highly costly and sensitive to variations in input resolution.
-To overcome these limitations, a training-free video editing approach that conditions on both image and text prompts is essential.
-However, a key challenge remains: maintaining temporal consistency while preserving motion dynamics is fundamental to video editing, but becomes particularly difficult in a training-free setting, leading to color shifts, shape distortions, or loss of fine details.
-As shown in Fig.
-2(b), simply integrating an image prompt into an existing T2V model without fine-tuning results in significant quality degradation, particularly in object shape fidelity.
-Figure 2: Importance of appropriate integration of image prompts in video editing.
-Given (a) a source video, we compare (b) a state-of-the-art text-driven video editing method, (c) a naive combination of a pretrained T2V model with an image-guided module, and (d) our proposed method.
-The text-only approach in (b) struggles to accurately transfer the fine-grained details of the car in the reference image.
-Simply attaching the image-guided module to a pretrained T2V model, as in (c), produces unnatural results.
-In contrast, (d) our proposed training-free method precisely transfers the car in the reference image into the video with high fidelity.
-Therefore, we propose that well-structured noise maps, generated through a controlled combination of the source video’s latent representation and Gaussian noise, effectively preserve motion consistency even without additional training.
-These noise maps act as a key component in achieving high-quality video edits.
-To further refine the noise maps to ensure seamless object transitions, we introduce a dilated dual masking strategy that blends the target object with the original video, which becomes particularly effective in handling shape discrepancies between the source and target objects.
-We also observe that a simple yet effective enhancement to the existing negative image prompt formulation used in IP-Adapter ipadapter effectively leads to realistic edits without requiring additional training.
-Building on these findings, we present ImEdit, the first zero-shot, training-free video editing approach conditioned on both image and text prompts, enabling high-fidelity object replacement while preserving temporal consistency.
-Our method leverages well-structured craft noise maps for motion preservation, dilated dual masking for seamless object replacement, and zero image guidance for color fidelity.
-Additionally, Our approach is highly flexible to varying image resolutions, by leveraging a T2V backbone pretrained on diverse resolutions without additional fine-tuning, ensuring robust generalization across different input scales.
+This challenge becomes even more pronounced when the editing method necessitates one-shot finetuning for each edit [videop2p, fatezero, save].
+To mitigate the need for costly text prompt engineering, recent work has explored video editing methods guided by both image and text.
+However, such approaches [genvideo, make_a_pro] suffer from weakened temporal consistency and require per-video finetuning, making them unsuitable for real-world editing scenarios.
+Other approaches employ additional guidance such as user-defined semantic keypoints alongside image and text prompts [videoanydoor, videoswap] to achieve more precise control.
+While effective for temporally coherent local editing guided by reference images, these methods require substantial user interaction and additional training, limiting their practicality.
+Therefore, we propose VIdeo editing with NOise control (VINO), the first zero-shot, training-free video editing framework conditioned on both image and text prompts.
 As illustrated in Fig.
-2(c), our proposed method effectively modifies video objects based on reference images and simple text prompts, eliminating the need for complex prompt engineering.
-We also evaluate the time efficiency of the proposed method against baseline methods in Fig.
-3.
-Across all tested input video sizes, our method consistently achieves the fastest runtime, demonstrating its advantage as a training-free solution for video editing.
-Notably, Fig.
-1 highlights that our method effectively addresses challenging scenarios, including small source objects (top), significant color differences between the source and target objects (bottom left), and dynamic motion in the source video (bottom right), while maintaining motion consistency throughout the edited results.
-In contrast, baseline methods often fail to accurately reflect the reference appearance under these settings.
-Figure 3: Comparison of runtime across different methods. Results are averaged over three runs for each unique input video size, with standard deviation.
-Through extensive quantitative and qualitative comparisons, we demonstrate that our approach outperforms existing state-of-the-art video editing methods in both visual quality and temporal stability.
-By eliminating the need for fine-tuning, our method is efficient, scalable, and widely applicable across various video editing tasks.
-In summary, our key contributions are as follows:
-  • We propose ImEdit, the first zero-shot, training-free video editing approach that leverages both text and image prompts to seamlessly replace objects while preserving original motion dynamics and temporal consistency.
-  • We introduce an effective noise map construction strategy to guide the denoising U-Net for coherent and accurate video edits.
-As key components, $\rho$-start sampling preserves motion patterns and spatial layout of the source video, while dilated dual masking ensures smooth transitions in appearance and shape.
-  • To further enhance color saturation and visual fidelity, we present zero image guidance, a controllable negative image prompt strategy that blends a zero image embedding into the existing negative image guidance.
-  • Comprehensive quantitative and qualitative evaluations demonstrate that our method outperforms existing state-of-the-art approaches.
-Figure 4: Overview of our proposed training-free approach.
-ImEdit constructs two levels of noise maps to guide the editing process.
-First, a rough noise map is constructed from the source video and denoised using a pretrained T2V model with image and text guidance, producing a rough edit.
-Next, a refined good noise map is constructed and denoised under the same conditions to produce the final video.
-In both stage, we denoise for $\rho$ steps by employing $\rho$-start sampling.
-The output more faithfully reflects the input prompts while preserving motion and scene structure.
+LABEL:fig:teaser, our method enables localized video edits guided by reference images, without any additional training or cumbersome user interaction.
+We leverage recent advances in the image domain to enable high-fidelity video editing without finetuning.
+In particular, pretrained image encoders and adapters [pbe, chen2024zero, stableviton, ipadapter] have shown success in injecting reference features into cross-attention layers.
+However, directly connecting these image-based techniques to existing text-to-video (T2V) models produces unnatural results, pasting the reference image onto the video without considering the source video’s layout (Fig.
+2(c)).
+We address this semantic unnaturalness by introducing $\rho$-start sampling inspired by [sdedit], which starts DDPM sampling from intermediate time steps to generate realistic images from stroke paintings.
+In our strategy, both DDIM sampling and inversion begin from intermediate time steps, preserving the layout and motion of the source video.
+While this strategy effectively balances the appearance of the reference and the layout of the source, it still struggles when replacing source objects with targets of significantly different sizes or shapes (Fig.
+7(a)).
+To overcome object size and shape discrepancies, we design VINO with a two-stage dilated dual masking strategy.
+In the first stage, we apply dilated masking to blend source video latents with editing latents during the sampling process, producing a coarse edited video.
+In the second stage, this coarse edited video is refined using dilated dual masking, enabling more flexible spatial blending for accurate appearance transfer.
+Figure 4: Overview of VINO.
+Our proposed training-free approach, VINO, adopts a coarse-to-fine scheme to guide the two-stage editing process via strategically designed noise maps.
+Both stages leverage $\rho$-start sampling with text and image prompts.
+In the first stage, a rough noise map is constructed from the source video via dilated masking (a-1), which is then denoised by the pretrained T2V model to yield a rough edit (b-1).
+In the second stage, a refined good noise map is derived from the rough edit via dilated dual masking (a-2).
+Finally, the model produces the final output based on this good noise map (b-2), faithfully reflecting the prompts while preserving motion and structure.
+We further present zero-image guidance to address occasional color desaturation.
+This controllable negative prompt strategy steers generation away from the zero-image direction, thereby improving color fidelity.
+Consequently, our method achieves high visual quality and temporal consistency (Fig.
+2(d)), while offering clear efficiency advantages over methods requiring finetuning (Fig.
+3).
+In summary, our key contributions are:
+  • We propose VINO, the first zero-shot video editing method using both text and image prompts, preserving motion and temporal consistency without training.
+  • We introduce a noise map construction strategy including $\rho$-start sampling to preserve the motion patterns and spatial layout of the source video, and dilated dual masking to ensure smooth appearance transfer.
+  • We present zero image guidance, enhancing color saturation via controllable negative prompt blending.
+  • Comprehensive quantitative and qualitative evaluations demonstrate that our method outperforms existing state-of-the-art approaches in both quality and efficiency.
 2 Methods
-We introduce ImEdit, a training-free video editing approach that performs object replacement while preserving the overall coherence of the scene, guided by text or image prompts.
+We introduce VINO, a training-free video editing approach that performs object replacement while maintaining overall scene coherence using text and image prompts, as illustrated in Fig.
+4.
 Our method is grounded in the careful integration of well-structured noise with appropriately chosen components.
 Specifically, we consider two key ingredients: the foreground latent, obtained via $\rho$-start sampling, and the background latent, acquired through DDIM inversion of the source video.
-These latents are blended using a carefully designed mask derived from both the source and target object masks, employing dilated dual masking to ensure smooth transitions and spatial consistency.
+These latents are blended using a tailored mask derived from both the source and target object masks, employing dilated dual masking to ensure smooth transitions and spatial consistency.
 In addition, we apply zero image guidance, a negative prompt-based strategy, to enhance color saturation and overall visual fidelity in the final output.
 2.1 Preliminaries
-Latent diffusion models (LDMs)
-In diffusion-based image generation models, LDMs have become a standard approach due to their computational efficiency and ability to generate high-resolution, photorealistic images dalle2 ; stablediffusion ; imagen .
-These models leverage a pretrained autoencoder vae , consisting of an encoder $\mathcal{E}(\cdot)$ and a decoder $\mathcal{D}(\cdot)$, to map input images into a lower-dimensional latent space.
-Given an input image ${\bm{x}}$, the clean latent variable is obtained as ${\bm{z}}_{0}=\mathcal{E}({\bm{x}})$.
-During the forward process, gaussian noise is gradually added to the latent variable over $T$ steps.
-In the reverse process, a denoising network $\epsilon(\cdot)$ predicts noise at each step to iteratively refine the latent variable.
-Once the denoising process is complete, the decoder $\mathcal{D}(\cdot)$ transforms the latent variable back into the image space to generate the final output.
-DDIM sampling and inversion
-[…]
-[…]
-DDIM inversion predicts the sequence of noisy latents $\{{\bm{z}}_{t}\}_{t=1}^{T}$ from a given clean sample ${\bm{z}}_{0}$, reconstructing the corresponding diffusion trajectory without requiring any model retraining ddim ; inverting .
-We denote this inversion trajectory as DDIM-inv$({\bm{z}}_{0})$.
+Latent Diffusion Models (LDMs).
+LDMs [stablediffusion] perform the diffusion process in a compressed latent space obtained from an autoencoder $\mathcal{E},\mathcal{D}$.
+Given an input image $\bm{x}$}, the encoder produces a clean latent representation ${\bm{z}}_{0}=\mathcal{E}({\bm{x}})$, where the diffusion process adds and removes noise as:
+${\bm{z}}_{t}=\sqrt{\bar{\alpha}_{t}}\,{\bm{z}}_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(0,{\mathbf{I}}).$ (1)
+To balance realism and prompt adherence, classifier-free guidance [CFG] combines conditional and unconditional diffusion model predictions as follows:
+$\hat{\bm{\epsilon}_{\theta}}=(1+w)\bm{\epsilon}_{\theta}({\bm{z}}_{t},t,{\bm{c}})-w\,\bm{\epsilon}_{\theta}({\bm{z}}_{t},t,\bar{{\bm{c}}}),$ (2)
+where $w$ is the guidance scale, $\bm{c}$} is the conditioning input, and $\bar{{\bm{c}}}$ denotes the negative conditioning input.
+This design enables efficient training and sampling while maintaining high-quality visual fidelity consistently.
+Figure 5: Controlled noise sampling for structured edits in VINO.
+Foreground and background latents, obtained via $\rho$-start sampling and zero image guidance, are blended using a final mask formed by the union of dilated source and target masks.
+This enables clean removal of the source object and seamless insertion of the target.
+DDIM Sampling and Inversion.
+DDIM [ddim] provide a deterministic and efficient sampling process that reduces the number of steps required for faithful generation. The sampling process updates the latent as:
+${\bm{z}}_{t-1}$ $=\sqrt{\alpha_{t-1}}\left(\frac{{\bm{z}}_{t}-\sqrt{1-\alpha_{t}}\,\bm{\epsilon}_{\theta}({\bm{z}}_{t},t)}{\sqrt{\alpha_{t}}}\right)$ (3)
+$\quad+\sqrt{1-\alpha_{t-1}}\cdot\bm{\epsilon}_{\theta}({\bm{z}}_{t},t),$
+where we denote this update rule as $\text{DDIM-samp}({\bm{z}}_{t},t)$.
+The inversion process can be obtained by reversing these sampling steps, enabling reconstruction and fine-grained editing in the latent space.
+For editing, specific regions can be modified by blending DDIM inversion latents from the source and ongoing edit using a spatial mask:
+${\bm{z}}^{\prime}_{t}={\bm{z}}^{\text{{fg}}}_{t}\odot{\bm{M}}+{\bm{z}}^{\text{{bg}}}_{t}\odot(1-{\bm{M}}),$ (4)
+where $\bm{M}$} specifies the edited region, ${\bm{z}}^{\text{{fg}}}_{t}$ and ${\bm{z}}^{\text{{bg}}}_{t}$ denote background and foreground latents, respectively.
 2.2 Overview
@@
 Similarly, $C_{r},H_{r},$ and $W_{r}$ represent the channel and spatial dimensions of the reference image.
-Our approach achieves this goal through a two-stage editing pipeline, as illustrated in Fig.
+Our approach achieves this goal through a two-stage editing pipeline, as shown in Fig.
 4.
@@
 We first encode the source video $\mathcal{V}_{s}$ into the latent space using a pretrained VAE encoder $\mathcal{E}$.
-We then initialize the starting noisy latent using $\rho$-start sampling (see Section 2.3.2), a modification of the standard DDIM sampling schedule, which enables better preservation of the source video’s structure and motion.
-This noisy latent is passed into a pretrained T2V diffusion model $\epsilon_{\theta}$ where we incorporate both image and text guidance using the IP-Adapter ipadapter .
-The image and text features from CLIP image encoder $\text{CLIP}_{\text{img}}$ and CLIP text encoders $\text{CLIP}_{\text{txt}}$ clip are injected into the each corresponding cross-attention layers of the T2V model as condition embeddings:
+We then initialize the noisy latent using $\rho$-start sampling (Section 2.3.2), a modification of the DDIM sampling schedule, which enables better preservation of the source video’s structure and motion.
+This noisy latent is passed into a pretrained T2V diffusion model $\epsilon_{\theta}$ where we incorporate both image and text guidance using the IP-Adapter [ipadapter].
+The image and text features from CLIP image encoder $\text{CLIP}_{\text{img}}$ and CLIP text encoders $\text{CLIP}_{\text{txt}}$ [clip] are injected into each corresponding cross-attention layers of the T2V model as condition embeddings:
 ${\bm{c}}_{\text{img}}=\text{CLIP}_{\text{img}}(\mathcal{R})\quad\text{ and }\quad{\bm{c}}_{\text{txt}}=\text{CLIP}_{\text{txt}}(\mathcal{T}),$ (5)
 In addition, negative guidance using $\bar{{\bm{c}}}_{\text{img}}$ and $\bar{{\bm{c}}}_{\text{txt}}$, which denote the negative image and text embeddings respectively, helps suppress irrelevant visual or textual features.
-During the denoising process with condition embeddings, we employ a foreground-background separation mechanism to perform object-centric modifications while preserving background fidelity, inspired by latent blending latentblending ; blendeddiffusion ; videoswap ; genvideo .
+During the denoising process with condition embeddings, we employ a foreground-background separation mechanism to perform object-centric modifications while preserving background fidelity, inspired by latent blending [latentblending, blendeddiffusion, videoswap, genvideo].
 Specifically, we split the noisy latent ${\bm{z}}_{t}$ into foreground and background regions using the object segmentation mask $\bm{M}$}.
@@
 Formally, the foreground latent ${\bm{z}}_{t}^{\text{{fg}}}$ and background latent ${\bm{z}}_{t}^{\text{{bg}}}$ at time step $t$ are computed as:
-[…]
-${\bm{z}}^{\text{{bg}}}_{t\texttt{-}1}$ $={\bm{z}}_{t\texttt{-}1},\ \text{with }{\bm{z}}_{t-1}\in\{{\bm{z}}_{t}\}_{t=1}^{\tau_{\nu}}\text{ from }\text{DDIM-inv}({\bm{z}}_{0}).$ (4)
-We then blend these two latents using ${\bm{M}}$ via our dilated dual masking (Section 2.3.1), as follows:
+${\bm{z}}^{\text{{fg}}}_{t\texttt{-}1}$ $=\text{DDIM-samp}({\bm{z}}^{\prime}_{t},{\bm{c}}_{\text{img}},{\bm{c}}_{\text{txt}},\bar{{\bm{c}}}_{\text{img}},\bar{{\bm{c}}}_{\text{txt}},t),$ (6)
+${\bm{z}}^{\text{{bg}}}_{t\texttt{-}1}$ $={\bm{z}}_{t\texttt{-}1},\ {\bm{z}}_{t-1}\in\{{\bm{z}}_{t}\}_{t=1}^{\tau_{\nu}}\text{ from }\text{DDIM-inv}({\bm{z}}_{0}).$ (7)
+We then blend these two latents using $\bm{M}$} via our dilated dual masking (see Section 2.3.1), as follows:
 ${\bm{z}}^{\prime}_{t\texttt{-}1}={\bm{z}}^{\text{{fg}}}_{t\texttt{-}1}\odot{\bm{M}}+{\bm{z}}^{\text{{bg}}}_{t\texttt{-}1}\odot(1-{\bm{M}}),$ (8)
@@
 Through the denoising process in the first stage, we obtain a roughly edited video $\mathcal{V}_{r}$, which reflects the general structure and appearance of the target object.
-However, this rough edit may still retain residual features of the source object, which should be removed for a clean transformation.
+However, this rough edit may still retain residual features of the source object, which should be removed for a clean visual transformation ultimately.
 To address the limitations of single mask blending, we introduce a second stage that further refines the output.
@@
 Note that the entire pipeline is training-free, relying solely on pretrained models for mask extraction and video generation.
-Figure 5: Effective noise construction using our dilated dual masking.
-We blend the foreground and background latents, obtained through our proposed methods such as $\rho$-start sampling and zero image guidance, using a final mask constructed by taking the union of the dilated source and target masks.
-This enables clean removal of the source object and seamless replacement with the target.
 2.3.1 Dilated dual masking
-The effectiveness of our method stems from the careful coordination of structured noise using the proposed dilated dual mask, which plays a crucial role in defining the editable regions within each frame.
-This mask is central to enabling selective editing of target objects, as illustrated in Fig.
-5.
-In the first stage, only the source object mask ${\bm{M}}_{s}$ is available, which roughly outlines the foreground region to be edited.
+Our proposed method coordinates structured noise to define editable regions in each frame and enables selective editing of target objects (Fig.
+5).
+Unlike Blended Latent Diffusion (BLD) [blendeddiffusion], which gradually shrinks masks during diffusion to address resolution degradation, our dilated dual masking uses a fixed dilation to refine boundaries and better spatially align source–target shapes.
+In the first stage, only the source object mask ${\bm{M}}_{s}$ is available, which roughly outlines the foreground region to be edited (Fig.
+4(b-1)).
 However, this mask lacks contextual adaptation to the background, often resulting in sharp edges and unnatural transitions at the object boundaries.
-Thus, we apply morphological dilation to smooth boundaries and facilitate seamless object replacement.
+Thus, we apply morphological dilation to smooth boundaries and facilitate seamless object replacement (Fig.
+4(a-1)).
 Specifically, a single-iteration morphological dilation is applied to a binary matrix $M$ with a kernel of size $k\times k$.
 For each pixel position $(p,q)$, the output value is determined as the maximum value within a neighborhood defined by ${\mathcal{N}}_{k}$, where ${\mathcal{N}}_{k}=\{(i,j)\mid-r\leq i,j\leq r\}$ and $r=\lfloor\frac{k}{2}\rfloor$.
-The dilation operation is formally expressed as follows:
+The dilation operation is formally expressed as:
 ${{\bm{M}}^{\prime}}[p,q]=\max_{(i,j)\in{\mathcal{N}}_{k}}{\bm{M}}[p+i,q+j].$ (9)
 As a result, the dilated source mask ${\bm{M}}^{\prime}_{s}$ is used to blend latents in Eq. 8 to produce rough edits $\mathcal{V}_{r}$.
-In the second stage, we similarly obtain the dilated rough mask ${\bm{M}}^{\prime}_{r}$ from the rough edits.
-However, blending foreground and background latents using either the source mask ${\bm{M}}^{\prime}_{s}$ or the rough mask ${\bm{M}}^{\prime}_{r}$ poses the risk that parts of the source video may remain visible or that the generation area for the target object may not be fully covered.
-To address this, we leverage masks from both the first and second stages.
-The final mask ${\bm{M}}_{e}$ is computed as the union of the source mask ${\bm{M}}^{\prime}_{s}$ and the rough mask from the first stage ${\bm{M}}^{\prime}_{r}$, as shown in the following equation:
+In the second stage, we similarly obtain the dilated rough mask ${\bm{M}}^{\prime}_{r}$ from the rough edits $\mathcal{V}_{s}$ (Fig.
+4(b-2)).
+However, blending foreground and background latents using either the dilated source mask ${\bm{M}}^{\prime}_{s}$ or the dilated rough mask ${\bm{M}}^{\prime}_{r}$ poses the risk that parts of the source video may remain visible or that the generation area for the target object may not be fully covered.
+To address this, we leverage masks from both the first and second stages (Fig.
+4(a-2)).
+The final mask ${\bm{M}}_{e}$ is computed as the union of the dilated source mask ${\bm{M}}^{\prime}_{s}$ and the dilated rough mask from the first stage ${\bm{M}}^{\prime}_{r}$, as shown in the following equation below:
 ${\bm{M}}_{e}={\bm{M}}^{\prime}_{s}\cup{\bm{M}}^{\prime}_{r}.$ (10)
 Finally, we construct the good noise ${\bm{z}}^{\prime}_{t}$ using the final mask ${\bm{M}}_{e}$ to guide the final blending process in Eq.
-5.
+8.
 This enables the second stage to generate more realistic, artifact-free results, as ${\bm{M}}_{e}$ precisely defines the edited region.
-In particular, it ensures clean removal of the original object, accurate reconstruction of the new object’s shape, and consistent preservation of the background.
+In particular, it ensures clean removal of the original object, accurate reflection of the new object’s shape, and consistent preservation of the background appearance consistently.
 2.3.2 $\mathbf{\rho}$-start sampling
-Preserving consistency with the source video is critical in video editing, particularly for maintaining visual structure and motion save ; camel ; deco .
-We hypothesize that initiating the sampling process from a less perturbed latent, closer to the source video, facilitates the preservation of these characteristics.
-To this end, we propose an alternative sampling strategy that modifies the standard DDIM schedule by starting the denoising process from a less noisy latent representation, thereby improving structural and temporal fidelity.
-In the conventional reduced DDIM schedule ddim , sampling proceeds along a trajectory ${{\bm{z}}_{\tau_{1}},\ldots,{\bm{z}}_{\tau_{\nu}}}$, starting from a highly noisy latent ${\bm{z}}_{\tau_{\nu}}$.
+Preserving consistency with the source video is critical in video editing, particularly for maintaining structural integrity and coherent motion [save, camel, deco].
+Inspired by SDEdit [sdedit], we hypothesize that initiating the sampling process from a less perturbed latent, closer to the source video, helps retain both spatial structure and temporal dynamics.
+To this end, we propose an alternative sampling strategy that modifies the standard DDIM schedule by starting the denoising process from a less noisy latent representation, thereby improving structural and temporal fidelity (Fig.
+4(d)).
+In the conventional reduced DDIM schedule [ddim], sampling proceeds along a trajectory ${\bm{z}}_{\tau_{1}},\ldots,{\bm{z}}_{\tau_{\nu}}$}, starting from a highly noisy latent ${\bm{z}}_{\tau_{\nu}}$.
 In contrast, our approach begins from an earlier timestep, ${\bm{z}}_{\tau_{\rho}}$, where $\rho<\nu$.
-This shortens the denoising trajectory to only $\rho$ steps, $\{{\bm{z}}_{\tau_{1}},\ldots,{\bm{z}}_{\tau_{\rho}}\}$, facilitating the retention of more information from the source video.
+This shortens the denoising trajectory to only $\rho$ steps, $\{{\bm{z}}_{\tau_{1}},\ldots,{\bm{z}}_{\tau_{\rho}}\}$, better preserving source content overall.
 The parameter $\rho$ controls how early the sampling begins in the reverse diffusion process.
@@
 By adjusting $\rho$, we gain fine-grained control over the degree of signal retention, enabling higher fidelity in both motion and overall scene consistency.
+Table 1: Quantitative comparison.
+We compare VINO against four state-of-the-art models and two naive adaptations of image editing and video generation models for video editing.
+We measure performance using both automatic metrics and human evaluations to assess text, image, and motion alignment, temporal consistency, and overall quality.
+Method Model support Model scoring metrics Human evaluations
+Text Image Train-free DINO $\uparrow$ CLIP-T $\uparrow$ Temp $\uparrow$ PSNR $\uparrow$ Image $\downarrow$ Text $\downarrow$ Motion $\downarrow$ Temp $\downarrow$ Overall $\downarrow$
+SDEdit-Video [sdedit] 0.431 0.303 0.959 16.3 4.89 5.08 5.93 5.02 5.36
+FateZero [fatezero] 0.438 0.311 0.967 20.1 4.59 4.53 4.44 5.06 4.75
+Video-P2P [videop2p] 0.424 0.305 0.971 19.0 4.00 4.07 3.68 4.09 3.70
+TokenFlow [tokenflow] 0.480 0.325 0.975 19.6 3.78 3.09 2.61 2.50 2.95
+AnimateDiff [animatediff]+IPA [ipadapter] 0.399 0.308 0.955 24.7 4.59 5.01 5.24 4.30 4.93
+Make-A-Pro [make_a_pro] 0.444 0.309 0.965 16.9 3.93 4.24 3.64 5.00 4.35
+[…]
+Figure 6: Qualitative comparison.
+We compare the generation quality of VINO against four state-of-the-art baseline models.
+VINO better preserves reference appearance across frames, while baselines struggle to retain key visual features of the target object.
 2.4 Zero image guidance
-To enhance color fidelity, we introduce zero image guidance, helping better align with the appearance of the reference image.
-Our refined guidance strategy blends the conventional negative image embeddings CFG ; glide , represented by a zero vector of the same size as the CLIP image embedding, with a zero image guidance, which refers to the CLIP image embedding derived from an all-zero image.
-The final negative image guidance $\bar{{\bm{c}}}_{\text{img}}$ is computed as a weighted sum of the two, controlled by a blending factor $\gamma$, and can be formally expressed as:
-$\bar{{\bm{c}}}_{\text{img}}=\gamma\cdot E_{I}({\mathbf{I}}_{0})+(1-\gamma)\cdot\mathbf{0},$ (8)
-where $\gamma\in[0,1]$ is a hyperparameter, $\mathbf{0}\in\mathbb{R}^{d}$ is a zero vector with the same size as the CLIP image embedding, $E_{I}(\cdot)$ represents the CLIP image encoder, and ${\mathbf{I}}_{0}\in\mathbb{R}^{C\times H\times W}$ is an all-zero image input.
-This formulation enables a tunable trade-off between the two types of negative guidance, allowing finer control over the color intensity of the generated object.
-The resulting $\bar{{\bm{c}}}_{\text{img}}$, together with the conventional $\bar{{\bm{c}}}_{\text{txt}}$, serves as negative guidance during the sampling process in Eq.
-3.
+To enhance color fidelity, we present zero image guidance (Fig.
+4(c)).
+It blends the conventional negative image embeddings [CFG, glide], defined by a zero vector of the same size as the CLIP image embedding, with a zero image guidance, which refers to the CLIP image embedding derived from an all-zero image.
+The final negative image guidance $\bar{{\bm{c}}}_{\text{img}}$ is computed as a weighted sum of the two, controlled by a blending factor $\gamma$:
+$\bar{{\bm{c}}}_{\text{img}}=\gamma\cdot\text{CLIP}_{\text{img}}({\mathbf{I}}_{0})+(1-\gamma)\cdot\mathbf{0},$ (11)
+where $\gamma\in[0,1]$ is a hyperparameter, $\mathbf{0}\in\mathbb{R}^{d}$ is a zero vector with the same size as the CLIP image embedding, $\text{CLIP}_{\text{img}}(\cdot)$ is the CLIP image encoder, and ${\mathbf{I}}_{0}\in\mathbb{R}^{C\times H\times W}$ is an all-zero image.
+This formulation offers a parameterized trade-off between the two types of negative guidance, allowing finer control over the color intensity of the generated object.
+The resulting $\bar{{\bm{c}}}_{\text{img}}$, combined with the conventional $\bar{{\bm{c}}}_{\text{txt}}$, is used as negative guidance during the sampling process in Eq.
+6.
+Figure 7: Ablation study on the dilated dual masking.
+The output of the first stage using the source video $\mathcal{V}_{s}$ and source object mask is shown in (a).
+Subsequently, the outputs of the second stage from the intermediate video $\mathcal{V}_{r}$ using different masking strategies are presented in (b)-(d).
+Masks are overlaid on the corresponding video.
+The proposed dilated dual masking (d) achieves the most accurate edits.
+Figure 8: Ablation study on the effect of varying the $\rho$ value.
 3 Experiments
-3.1 Experimental setup
-Evaluation metrics.
-We conduct evaluations on 18 pairs of video and reference image, using videos collected from DAVIS dataset davis and the dataset released by Video-P2P videop2p .
-To assess how well the generated videos align with the given text and image prompt, we use CLIP-T score and DINO score.
-Specifically, CLIP-T score measures the alignment between the target text prompt and the generated videos using CLIP text encoder clip , while DINO score calculates the similarity between the reference image and the generated video using DINO dino .
-To evaluate temporal consistency, we extract CLIP image embeddings for each frame in the edited video using the CLIP image encoder clip and measure the cosine similarity between embeddings of adjacent frames.
+Dataset.
+We evaluate VINO on 20 video-image pairs.
+Videos are sourced from DAVIS [davis], Video-P2P [videop2p], and a curated set covering various object categories (e.g., humans, animals, everyday items) to ensure broad semantic and visual diversity.
+Reference images include both real and synthetic samples generated by models such as GPT-4o and Gemini.
 Baseline models.
-We compare our proposed method with two baselines: Video-P2P videop2p and Make-A-Protagonist (Make-A-Pro) make_a_pro .
-To the best of our knowledge, Make-A-Pro is the only baseline that directly aligns with our problem setting and offers publicly released code among existing methods.
-However, in order to comprehensively demonstrate the effectiveness of our image prompt approach, we also include Video-P2P, a state-of-the-art text-driven video editing model, as an additional baseline.
+We compare VINO against six representative video editing methods, as summarized in Table 1, which use text or image prompts for semantic appearance control.
+SDEdit-Video [^2] We use the HuggingFace implementation of SDEdit extended to video.
+[sdedit], FateZero [fatezero], Video-P2P [videop2p], and TokenFlow [tokenflow] represent text-guided video editing methods.
+To our knowledge, Make-A-Protagonist (Make-A-Pro) [make_a_pro] is the only publicly available method that supports both image and text conditioning, making it the most directly comparable to VINO.
+AnimateDiff [animatediff] +IP-Adapter (IPA) [ipadapter] serves as a modular baseline that integrates a pretrained T2V model with an image-guided adapter, illustrating the limitations of naive fusion strategies.
 Implementation details.
-We use AnimateDiff animatediff as the base T2V model, IP-Adapter ipadapter v1.5 for the reference image encoding, and SAM2 sam2 for obtaining masks.
-We conduct experiments with videos of various resolutions (320$\times$576, 480$\times$480, 512$\times$512, 600$\times$600 and 640$\times$640) and different numbers of frames (8 and 16).
-For text prompts, we use the same descriptions when comparing our method with Video-P2P videop2p .
-In contrast, for Make-A-Pro make_a_pro , we follow their strategy by extracting protagonist information using BLIP-2 blip2 and generating prompts via visual question answering.
-Additionally, since Video-P2P videop2p only supports square outputs, we paste its generated video back into the corresponding regions of the source video to enable a comparison.
-All experiments are conducted on a single NVIDIA A100 GPU with 40GB memory.
-3.2 Experimental results
-3.2.1 Comparisons with state-of-the-art methods
-Quantitative and qualitative results.
-We quantitatively evaluate our method with a focus on alignment with text and image prompts, visual feature similarity, and temporal consistency across video frames.
-As shown in Table 1, our method consistently outperforms all baselines across all evaluation metrics, both automatic and human.
-In terms of model scoring metrics, our method achieves the highest CLIP-T, DINO, and Temp scores, indicating superior text alignment, object similarity, and temporal consistency, respectively.
-These trends are further supported by human evaluation.
-Our method receives the lowest scores in Image, Text, and Temp categories, demonstrating that it generates objects most similar to the target, aligns most faithfully with the input text prompts, and maintains the most consistent appearance across frames.
-In addition, our method is rated highest in Overall quality, clearly confirming its effectiveness and perceptual superiority over the baselines.
-Qualitative comparisons further demonstrate the effectiveness of our method in accurately preserving both motion and identity across diverse editing scenarios, outperforming the baselines.
-In the left example of Fig.
-6, Make-A-Pro make_a_pro preserves the motion from the source video but fails to accurately reflect the identity of the reference image, even struggling to edit the dog into a fox.
-Video-P2P videop2p partially captures the identity of the fox but fails to preserve motion, resulting in an incorrect number of objects, with two foxes appearing in a single frame.
-In contrast, our method produces the most accurate edited video, successfully maintaining both the motion from the source video and the identity of the reference object.
-The right example in Fig.
-6 features a source video with less dynamic movement.
-In this case, all methods preserve motion, but only our method correctly generates a white duck as specified in both the text condition and reference image.
-This example highlights the ability of our method to effectively handle both image and text prompts in video editing tasks.
-Table 1: Quantitative comparison results.
-Quantitative metrics include CLIP-Text similarity (CLIP-T), visual feature similarity (DINO), and temporal consistency (Temp).
-Human evaluations measure the text-video alignment (Text), visual fidelity to the target object (Image), motion consistency with the source video (Motion), temporal consistency (Temp), and overall visual quality (Overall).
-Method Model scoring metrics Human evaluations
-CLIP-T $\uparrow$ DINO $\uparrow$ Temp $\uparrow$ Text $\downarrow$ Image $\downarrow$ Motion $\downarrow$ Temp $\downarrow$ Overall $\downarrow$
-Video-p2p videop2p 0.300 0.393 0.961 2.01 2.05 2.27 1.99 2.17
-Make-A-Pro make_a_pro 0.294 0.362 0.958 2.66 2.61 1.99 2.64 2.57
-ImEdit (Ours) 0.326 0.474 0.966 1.33 1.34 1.74 1.37 1.26
-Figure 6: Qualitative comparison results on DAVIS dataset.
-We compare the generation quality of our method against baseline methods.
-While the baseline methods fail to incorporate key visual features from the reference image, our method effectively transfers the target object into the source video, faithfully preserving its appearance across frames.
+Our main experiments adopt AnimateDiff [animatediff] as the base T2V model, IP-Adapter v1.5 [ipadapter] for reference image conditioning, and SAM2 [sam2] for mask generation.
+All experiments are conducted on a single NVIDIA A100 GPU.
+Additional implementation details, including hyperparameters, baseline setups, and extensions to other T2V backbones, are provided in Appendix E.1.
+Due to space constraints, we present representative qualitative comparisons and ablation results in the main paper, with full qualitative results provided in Appendix B.
+Figure 9: Ablation study on the image guidance scale $\gamma$.
+3.1 Comparisons with state-of-the-art methods
+Quantitative results.
+We evaluate VINO using four metrics: DINO score measures visual alignment via cosine similarity between features of the reference image and generated video [dino].
+CLIP-Text score quantifies text alignment by computing cosine similarity between CLIP embeddings of the generated video and the text prompt [clip].
+Temporal score assesses motion consistency through cosine similarity between CLIP image embeddings of adjacent frames [clip].
+PSNR score measures how accurately the background of the source video is preserved by computing pixel-level similarity between the generated and source frames.
+VINO achieves the highest DINO score among all baselines, indicating strong visual fidelity to the reference appearance.
+In CLIP-Text and temporal scores, it performs comparably to TokenFlow while outperforming other methods.
+Notably, VINO is compatible with existing methods such as TokenFlow and can serve as an additional appearance control module.
+This composition yields further performance gains, as detailed in the Appendix A.
+For PSNR, AnimateDiff+IPA and VINO noticeably outperform the other methods.
+The small gap between them, despite both using the same latent blending strategy, arises from the edited area.
+AnimateDiff+IPA modifies a comparatively smaller region, whereas VINO employs dilated dual masking, which introduces a larger generation area, leading to a marginally lower PSNR.
 User study.
-We conduct a user study based on five evaluation criteria: (1) image alignment, which measures how well the object in the generated video resembles the one in the reference image; (2) text alignment, which assesses how accurately the generated video matches the given text condition; (3) motion alignment, which evaluates how well the motion in the generated video aligns with that of the source video; (4) temporal consistency, which examines the consistency of the object and background across frames; and (5) overall quality, which considers all the aforementioned factors to determine how natural and realistic the generated video appears.
-We ask 32 participants to evaluate 9 comparisons each, ranking the given results based on these criteria.
-We compute the average rankings, and Table 1 shows that our method ranks highest across all criteria.
-3.2.2 Ablation study
-Quantitative results.
-We quantitatively evaluate the contribution of each component in our method, as shown in Table 2.
-Starting from the base model (a), the integration of zero image guidance (b) leads to an overall performance improvement.
-Incorporating $\rho$-start sampling (c) brings a slight increase in CLIP-Text alignment, while the DINO and Temporal scores slightly decrease, indicating that adjusting the starting sampling point alone is insufficient to resolve the trade-off between motion preservation and appearance fidelity.
-The inclusion of dual masking (d) contributes to a more substantial gain across all metrics.
-Finally, our proposed method with dilation (e) achieves the best performance, with notable improvements.
-These results highlight the effectiveness of the complete method.
+We conduct a user study with 25 participants to assess perceptual quality across five criteria: image alignment, text alignment, motion alignment, temporal consistency, and overall quality.
+VINO consistently outperforms all baselines across all criteria, demonstrating strong perceptual fidelity in both appearance and motion.
+Qualitative results.
+Qualitative comparisons, shown in Fig.
+6, further demonstrate the effectiveness of VINO in preserving object appearance and motion across various editing scenarios, consistently outperforming baselines.
+In the man-to-panda translation, baselines either fail to generate a panda or produce artifacts or distortions, while VINO yields a clean and consistent result.
+A similar trend appears in the duck-to-chicken translation, where most methods generate generic chicken-like color rather than matching the reference, whereas VINO preserves its specific visual traits.
+In the rabbit-to-fox translation, most baselines struggle to maintain appearance fidelity or temporal coherence.
+Make-A-Pro, in particular, suffers from temporal flickering, including disappearing or reappearing limbs.
+In contrast, VINO maintains the correct fox appearance while preserving the background, producing coherent structure and stable motion throughout the video.
+Table 2: Comprehensive ablation study of VINO.
+We evaluate the contribution of each component to the overall performance of our method: (a) base model, (b) $\rho$-start sampling, (c) dual masking, (d) dilation, and (e) zero image guidance.
+Method DINO $\uparrow$ CLIP-T $\uparrow$ Temp $\uparrow$
+(a) 0.399 0.308 0.955
+(a) + (b) 0.429 0.311 0.953
+(a) + (b) + (c) 0.466 0.316 0.954
+(a) + (b) + (c) + (d) 0.474 0.316 0.955
+[…]
+3.2 Ablation study
 Effectiveness of dilated dual masking.
-We ablate our proposed dilated dual masking by progressively incorporating each masking schemes, as shown in Fig.
-7.
-We begin with the simplest configuration: a single stage editing setup that relies solely on the source mask.
-When using the source mask only, the generated object is often forced to conform to the contours of the source object, leading to unrealistic shapes, limited fidelity, and artifacts.
-This issue becomes particularly pronounced when there is a significant shape difference between the source and target objects.
-As shown in Fig.
-7(a), the source object (grey car) is structurally different from the target object; it is shorter in length, taller in height, and more angular than the target object, but these differences are not considered.
-Adding an additional stage to generate the rough mask and using only this mask alleviates the problem of the imprecise reflection of the target object’s features.
-However, Fig.
-7(b) reveals that it fails to completely remove the source, as the top part of the grey car still remains.
-To further address this, we take the union of the source and rough mask.
-While this reduces residuals from the source object, artifacts persist along the boundaries.
-Furthermore, fine-grained details of the generated target object are lost due to the strict constraints imposed by the merged mask region, as illustrated in Fig.
-7(c).
-Our proposed dilated dual masking effectively addresses the aforementioned issues by applying dilation to the masks and this is evident in Fig.
-7(d).
-It allows the generated target object to extend beyond the strict union region, enabling natural shape adaptation and smoother blending into the surrounding context.
-As a result, the proposed dilated dual masking achieves both structural integrity and visual realism, demonstrating its superiority.
-Table 2: Comprehensive ablation study of ImEdit.
-We evaluate the contribution of each component to the overall performance of our method.
-(a) Base model.
-(b) Zero image guidance.
-(c) $\rho$-start sampling.
-(d) Dual masking.
-(e) Dilation.
-Method CLIP-Text $\uparrow$ DINO $\uparrow$ Temp $\uparrow$
-(a) 0.302 0.361 0.958
-(a) + (b) 0.304 0.405 0.960
-(a) + (b) + (c) 0.309 0.401 0.955
-(a) + (b) + (c) + (d) 0.316 0.435 0.958
-(a) + (b) + (c) + (d) + (e) (Ours) 0.326 0.474 0.966
-Figure 7: Ablation study on the proposed dilated dual masking.
-(a) shows the output of the first stage using the source video $\mathcal{V}_{s}$ and source object mask, while (b)-(d) present outputs of the second stage from the intermediate video $\mathcal{V}_{r}$ using different masking strategies.
-Masks are overlaid on the corresponding video.
-The proposed dilated dual masking (d) achieves the most accurate edits, underscoring its effectiveness for the object replacement.
-Figure 8: Effect of varying the $\rho$ value.
-The rightmost column shows the final output after the second stage, while all other columns, excluding the reference image and source video, show the results of the first stage under different values of $\rho$.
-By adjusting sampling start point, we balance how strongly the appearance of the target object is retained versus the source motion.
-Lower $\rho$ emphasizes the owl’s visual traits, while higher $\rho$  increasingly preserves the parrot’s original movement and posture.
+We conduct an ablation study on our dilated dual masking by progressively applying different masking schemes (Fig.
+7).
+Using only the source mask constrains the generation to the contours of the source object, often resulting in unrealistic shapes and artifacts when the source and target differ significantly (Fig.
+7(a)).
+Applying only the rough mask alleviates this issue but leaves visible remnants of the source object (Fig.
+7(b)).
+Taking the union of both masks suppresses these remnants but introduces boundary artifacts and limits the detail due to the rigid mask constraint (Fig.
+7(c)).
+In contrast, our dilated dual masking effectively mitigates these limitations by dilating both masks before fusion, enabling the generated object to extend beyond the strict union region.
+This flexibility promotes more natural shape adaptation and smoother blending with the surrounding context (Fig.
+7(d)).
+As a result, our dilated dual masking achieves both structural coherence and visual realism.
 Effectiveness of $\rho$-start sampling.
 As shown in Fig.
-8, varying the parameter $\rho$, which determines the sampling start point, enables control over the balance between the target appearance and source motion.
-Using a larger $\rho$ leads the model to fails to accurately follow the motion and spatial layout of the source video.
-For example, while the parrot in the source video is facing left, the owl in the edited video is facing forward or at an incorrect angle.
-In contrast, using a smaller $\rho$ allows the model to better capture the motion and spatial dynamics of the source video, resulting in the owl correctly facing sideways, consistent with the source object.
-However, if $\rho$ is too small, the transformation may become too weak, leading to incomplete replacement of the source object.
-With an appropriate choice of $\rho$  the edited video successfully balances appearance fidelity and motion preservation, highlighting the importance of temporal alignment in video editing tasks.
-Effectiveness of the zero image guidance.
-We analyze the impact of $\gamma$ in the proposed zero image guidance.
+8, adjusting $\rho$, which controls the sampling start point, enables a trade-off between target appearance and source motion.
+Larger $\rho$ values cause the model to deviate from the source motion and layout.
+For example, the parrot facing left in the source video becomes an owl facing forward in the edited result.
+Smaller $\rho$ values better preserve the source motion and spatial dynamics, producing an owl that correctly faces sideways.
+However, overly small $\rho$ weakens the transformation, leading to incomplete replacement of the source object.
+An appropriate choice of $\rho$ achieves a balanced edit, maintaining both appearance fidelity and motion preservation.
+Effectiveness of zero image guidance.
+We analyze the impact of $\gamma$ in our zero image guidance.
 As shown in Fig.
-9, increasing $\gamma$, which scales the contribution of the zero image guidance (see Eq.
+9, increasing $\gamma$, which scales the strength of the zero image guidance (see Eq.
 (11)), leads to more saturated colors.
-This suggests that mixing the zero image guidance into the negative image prompt effectively enhances color intensity, providing a simple yet powerful control for visual intensity.
-Figure 9: Effect of varying the $\gamma$ value. Increasing $\gamma$ enhances the influence of the zero image embedding, leading to more saturated color reproduction in the edited video.
+This suggests that incorporating the zero image guidance into the negative image effectively enhances color intensity, providing a simple yet powerful control for visual appearance.
+Quantitative results.
+Table 2 summarizes the effect of each component in our method.
+As we progressively add components (a)–(e), DINO and CLIP-Text scores increase monotonically, indicating that the proposed modules jointly improve appearance transfer from the reference image as well as adherence to the text prompt.
+In contrast, the temporal score remains relatively stable for variants (a)–(d), but improves significantly once zero image guidance (e) is introduced, leading our complete model to achieve the best performance across all metrics.
+These results quantitatively demonstrate that each component contributes effectively to temporally coherent video editing conditioned on both image and text.
 4 Conclusion
-In this work, we propose ImEdit, a novel zero-shot, training-free video editing method guided by both image and text prompts.
-ImEdit enables accurate object replacement with strong temporal coherence and achieves superior performance in both visual quality and motion consistency, outperforming existing state-of-the-art approaches with high efficiency and scalability.
+In this work, we propose VINO, a novel zero-shot, training-free video editing method guided by both image and text prompts.
+It enables accurate object replacement with strong temporal coherence, outperforming state-of-the-art methods in visual quality and motion consistency, and efficiency.
 Appendix
-This supplementary material complements the main manuscript, providing additional experimental results and implementation details.
-  • Appendix A provides an analysis of the effect of the DDIM inversion trajectory length.
-  • Appendix B provides an additional quantitative comparison with baseline methods.
+This supplementary material complements the main manuscript, providing additional experimental results and implementation details that could not be included because of space limitations.
+  • Appendix A illustrates the applicability of our method across various text-to-video models.
+  • Appendix B includes complete qualitative comparison and ablation studies.
   • Appendix C provides related work.
-  • Appendix D addresses the limitations of our method.
-  • Appendix E provides additional ablation results on dilated dual masking.
-  • Appendix F provides additional implementation details and user study details.
-  • Appendix G presents the overall algorithm of the proposed method.
-  • Appendix H provides additional qualitative results.
-Appendix A Effect of the DDIM inversion trajectory length
-In our latent blending process, the number of DDIM inversion sampling steps is treated as a hyperparameter with two discrete settings, 999 or 50, depending on the desired outcome.
-Using 999 steps often results in less vivid colors and a relatively noisy background.
-However, the edited object tends to exhibit more natural shapes and the overall video appears more realistic.
-In contrast, using 50 steps yields cleaner, more vivid appearances for both the object and background, and significantly reduces DDIM inversion time.
-Nonetheless, the shape of the edited object may appear less natural.
+  • Appendix D discusses the limitations of our method.
+  • Appendix E details additional experimental settings.
+  • Appendix F presents the overall algorithm of our method.
+  • Appendix G showcases additional qualitative results.
+Appendix A Applicability to various T2V models
+As shown in Table 3, our method is compatible with various text-to-video (T2V) models, enabling training-free video editing conditioned on both images and text.
+We evaluate the compatibility of our approach by extending the text-based generative model AnimateDiff [animatediff] and the text-based editing model TokenFlow [tokenflow] with our proposed components, including dilated dual masking, $\rho$-start sampling, zero-image guidance, and the image-guided module.
+Integrating our method into AnimateDiff yields higher scores across most metrics, demonstrating substantial improvements in structural consistency and visual similarity to the reference image.
+When applied to TokenFlow, DINO and PSNR also improve, indicating the integration strengthens reference alignment even in the text-based editing models.
+However, our method shows a lower temporal consistency score than that of text-prompt-only baselines.
+This reflects a trade-off inherent to image-based conditioning, where aligning each frame to a high-fidelity reference image introduces larger pixel-level variations, even though the edits remain perceptually coherent.
+Despite this trade-off, the consistent gains achieved with only marginal drops in temporal consistency across both generation and editing models underscore the applicability and generality of our method.
+Figure 10: Full qualitative comparison with baselines.
+We provide an extended version of Fig.
+6 from the main manuscript by including results from SDEdit-Video [sdedit] and AnimateDiff [animatediff]+IP-Adapter [ipadapter].
+Figure 11: Full qualitative results of the ablation study on dilated dual masking.
+Additional frames corresponding to Fig.
+7 in the main manuscript are provided to further illustrate the consistent transformation of the object across time.
+Appendix B Full qualitative comparison and ablations
+This section presents the full qualitative comparisons with baseline methods and qualitative ablation results that complement those in the main paper.
+Appendix B.1 includes complete qualitative comparison against baseline approaches, Appendix B.2 provides full results for ablation studies.
+Table 3: Quantitative evaluation of applicability to various T2V models.
+We integrate our proposed components into AnimateDiff and TokenFlow.
+The integration improves reference alignment and image quality with minimal trade-offs across different architectures.
+Method Model scoring metrics
+DINO $\uparrow$ CLIP-T $\uparrow$ Temp $\uparrow$ PSNR $\uparrow$
+AnimateDiff [animatediff] 0.412 0.308 0.980 13.9
++ Ours 0.498 0.321 0.972 24.7
+TokenFlow [tokenflow] 0.480 0.325 0.975 19.6
++ Ours 0.523 0.320 0.974 23.9
+B.1 Qualitative comparison with baselines
 As shown in Fig.
-10, the result with 999 steps shows a polar bear with reduced color saturation and increased background noise while the head and front paws are generated with a more natural structure.
-Conversely, the result with 50 steps shows a polar bear with more vivid white appearance and cleaner background, but the shape of the head and paws is less natural.
-Figure 10: Effect of the reduced inversion trajectory length. We evaluate how varying $\nu$, the number of DDIM steps used in inversion, affects editing quality.
-Appendix B Additional quantitative results
-Table 3 presents an additional quantitative comparison between ImEdit and existing baseline methods using CLIP-T, DINO, Temp, and PSNR, which respectively assess text-image alignment, the similarity between the reference image and the edited video, temporal consistency, and background preservation.
-To facilitate a more comprehensive evaluation, we extend the comparison in Table 1 by including FateZero fatezero and a naive variant of our approach, AnimateDiff animatediff combined with IP-Adapter ipadapter , and by adding PSNR as an additional metric.
-Specifically, AnimateDiff + IP-Adapter refers to the direct combining of our backbone model AnimateDiff with IP-Adapter, without any of the proposed components, as illustrated in Fig.
-2.
-FateZero fatezero is included as a baseline due to its strong editing performance among text-driven video editing methods that operate in a training-free setting, consistent with the setting of our method.
-Table 3: Additional quantitative comparison with baseline models.
-We compare our method (ImEdit) against existing baselines across four metrics: CLIP-Text similarity (CLIP-T), visual feature similarity (DINO), temporal consistency (Temp), and background preservation (PSNR).
-Method Model scoring metrics
-CLIP-T $\uparrow$ DINO $\uparrow$ Temp $\uparrow$ PSNR $\uparrow$
-FateZero fatezero 0.317 0.460 0.960 21.32
-Video-p2p videop2p 0.300 0.393 0.961 21.75
-Make-a-Pro make_a_pro 0.294 0.362 0.958 15.40
-AnimateDiff animatediff + IP-Adapter ipadapter 0.302 0.361 0.958 24.83
-ImEdit (Ours) 0.326 0.474 0.966 23.59
-As shown in Table 3, ImEdit outperforms Video-P2P videop2p , Make-a-Pro make_a_pro , and FateZero fatezero across all metrics.
-Notably, in terms of DINO scores, our approach achieves higher reference similarity not only compared to text-based methods such as Video-P2 videop2p and FateZero fatezero but also surpasses image-conditioned methods like Make-a-Pro make_a_pro .
-While AnimateDiff + IP achieves a higher PSNR than ours, this difference stems from the smaller edited region in its outputs.
-ImEdit employs dilated dual masking during editing, which results in a broader generation area and a slightly lower PSNR.
-Appendix C Related work
-C.1 Image and Video Generation
-Recent advancements in latent diffusion models (LDMs) have significantly improved the fidelity and controllability of image generation.
-Text-to-image (T2I) models such as Stable Diffusion stablediffusion , DALL-E 2 dalle2 , and Imagen imagen achieve high-resolution image synthesis aligned with complex text prompts.
-These models operate in the latent space for computational efficiency and support diverse conditioning mechanisms, including text, layout, and edge maps ipadapter ; controlnet .
-Their flexibility has made LDMs foundational components for extending generative modeling into the video domain.
-Motivated by T2I success, text-to-video (T2V) models have emerged to generate coherent video sequences conditioned on textual descriptions.
-CogVideo cogvideo adapts T2I backbones with temporal attention, enabling efficient video generation with minimal overhead.
-ZeroScope zeroscope , built on ModelScope modelscope , improves visual fidelity and consistency, producing higher-resolution outputs while mitigating common artifacts in earlier methods.
-AnimateDiff animatediff integrates temporal modules into Stable Diffusion stablediffusion , facilitating the T2V model with improved coherence between frames.
-Alongside architectural advances, controllability and identity preservation have gained importance.
-Several methods leverage a reference frame along with a prompt to guide spatial appearance and layout across time stiv ; ti2v ; controlvideo .
-In addition, others introduce subject-driven personalization via instance finetuning videoalchemist ; dreamvideo ; tune_a_video ; customcrafter , enabling video generation with identity consistency across diverse prompts.
-These trends illustrate a broader shift toward multimodal conditioning across text, image, and motion to allow more expressive and controllable video synthesis.
-C.2 Image and Video Editing
-Following the advances in image and video generation models, diffusion models have also made significant progress in image and video editing, allowing intuitive and precise content manipulation.
-In the image domain, Prompt-to-Prompt prompt-to-prompt introduces cross-attention manipulation within pretrained diffusion models, enabling semantic edits without finetuning or masks.
-InstructPix2Pix instructpix2pix leverages pairs of instructions and images to support editing guided by natural language.
-EditAnything editanything generalizes this paradigm by integrating multi-modal inputs such as text, masks, and user clicks.
-Video editing extends these capabilities to the temporal dimension.
-Tune-A-Video tune_a_video employs one-shot tuning of pretrained image diffusion models for video editing tasks, allowing users to modify video content and style with minimal data and strong temporal consistency.
-Video-P2P videop2p extends cross-attention control techniques from image to video, maintaining both spatial and temporal coherence.
-Text2Video-Zero text2video removes the need for paired text-video data, instead leveraging pretrained text-image models for flexible and data-efficient video editing.
-To overcome the limitations of editing based solely on text, recent studies propose image-conditioned video editing methods that use a reference image to guide appearance and identity genvideo ; make_a_pro ; cutandpaste .
-Make-A-Protagonist make_a_pro introduces a unified framework for protagonist replacement, background editing, and style transfer using mask-guided denoising.
-GenVideo genvideo proposes shape-aware masking and latent noise correction to ensure temporal consistency and subject fidelity.
-VideoSwap videoswap adds semantic point correspondences to provide interactive and precise control over the subject’s motion and appearance.
-Despite these advancements, existing methods still require a finetuning stage with the source video, which limits their overall effectiveness.
-In contrast, our approach eliminates this need entirely, enabling high-quality, zero-shot video editing with strong temporal and identity coherence.
-Appendix D Limitation
-Although our proposed method exhibits notable improvements over existing baselines, as evidenced by extensive experiments, it still faces certain limitations.
-When there is a substantial disparity in the innate physical characteristics between the source and target objects, our method struggles to perform object replacement.
-Figure 11 shows that our method generates wings for the red panda which is an anatomically implausible feature to transfer the parrot’s flying motion.
-Conversely, in attempting to preserve the red panda’s natural characteristics, the method fails to convey the intended motion.
-Figure 11: Example of failure case.
-A large semantic gap between the object in the source video and the reference image poses challenges for our method.
-We show results from two versions of our method, obtained by varying the sampling start point $\rho$ and the dilation value.
-In this case, editing a bird into a red panda results in missing or incorrect transfer of the wings and arms.
-Figure 12: Additional results on the proposed dilated dual masking.
-(a) shows the output of the first stage using the source video $\mathcal{V}_{s}$ and source object mask, while (b)-(d) present outputs of the second stage from the intermediate video $\mathcal{V}_{r}$ using different masking strategies.
-Masks are overlaid on the corresponding video.
-The proposed dilated dual masking (d) achieves the most accurate edits, underscoring its effectiveness for the object replacement.
-Appendix E Additional ablation results on dilated dual masking
-The proposed dilated dual masking enables the seamless replacement of the source object with the target object from the reference image across all frames.
+10, since SDEdit-Video [sdedit] and AnimateDiff [animatediff]+IP-Adapter [ipadapter] are adapted or composed from methods not originally designed for video editing, their qualitative results exhibit reduced visual fidelity compared to other baseline methods.
+Specifically, SDEdit-Video is extended from its original image-based formulation to the video domain, while AnimateDiff+IP-Adapter combines AnimateDiff with IP-Adapter in a naive manner.
+As a result, SDEdit-Video preserves only the coarse spatial layout of the source video and often introduces excessive noise, leading to degraded visual quality.
+In the case of AnimateDiff+IP-Adapter, reference images are directly pasted into the edited regions without sufficient consideration of semantic structure, resulting in visually unnatural artifacts.
+B.2 Qualitative ablation studies
+Dilated dual masking.
+Dilated dual masking enables the seamless replacement of the source object with the target object in the reference image across frames.
 Fig.
@@
 In Fig.
-12(a), when only the source mask is used, the target object is unnaturally constrained to the shape of the source mask, resulting in distorted generations.
+11(a), when only the source mask is used, the target object is unnaturally constrained to the shape of the source mask, resulting in distortion.
 In Fig.
-12(b), using only the rough mask fails to fully cover the source object, leaving undesired remnants of it in the edited video.
+11(b), using only the rough mask fails to fully cover the source object, leaving undesired remnants of it in the edited video.
 In Fig.
-12(c), employing a dual mask obtained by taking the union of the source and target masks alleviates artifacts from the source object and shape mismatches of the target object, but the mask boundaries still lack smoothness.
-In contrast, our proposed dilated dual mask in Fig.
-12(d) effectively mitigates these issues, enabling natural removal of the source object and coherent synthesis of the target object throughout the video.
-Appendix F Additional experimental settings
-Appendix F.1 provides additional implementation details, including hyperparameter settings. Appendix F.2 describes the evaluation questions used in our user study.
-F.1 Additional implementation details
-Our method involves four key hyperparameters that control different aspects of the video editing process, which include $\rho,\gamma,\text{dilation kernel size $k$, and guidance scale}$:
-  • $\rho\in[10,30]$: Control the starting timestep for the sampling process in $\rho$-start sampling, determining the trade-off between motion preservation from the source video and appearance transformation toward the reference image.
-  • $\gamma\in[0,1]$: Control the weighted combination of standard zero-vector and zero-image embeddings in zero image guidance, enabling fine-grained adjustment of color fidelity in the generated object.
-  • dilation kernel size $k$ $\in[0,10]$: Specifies the size of the dilation kernel used in the proposed dilated dual masking. This affects the spatial smoothness and coverage of the edited region.
-  • guidance scale $\in[6,10]$: Determine the strength of classifier-free guidance during denoising. A higher guidance scale enhances fidelity to the conditioning signal.
-F.2 User study details
-In the user study described in Section  3.2.1, participants are asked to evaluate five key aspects of the generated videos using the following ranking-based questions:
-  • Image Alignment: Rank the videos from 1st to 4th based on how well the appearance of the "object in the generated video" matches the "object in the reference image (excluding the background)".
-  • Text Alignment: Rank the videos from 1st to 4th based on how well the "generated video content" aligns with the "provided text prompt".
+11(c), employing a dual mask obtained by taking the union of the source and target masks, alleviates artifacts from the source object and shape mismatches of the target object, but the mask boundaries still lack smoothness.
+In contrast, our dilated dual mask in Fig.
+11(d) effectively mitigates these, enabling clean removal of the source object and coherent synthesis of the target object throughout the video.
+$\rho$-start sampling.
+Our $\rho$-start sampling initiates DDIM sampling and inversion from intermediate time steps, thereby balancing the appearance of the reference image with the spatial layout of the source video.
+Fig.
+12 supplements Fig.
+8 by additionally presenting the final edited video results.
+Specifically, while Fig.
+8 only illustrates the first stage edits under varying values of $\rho$, the rightmost column in Fig.
+12 provides the corresponding final outputs after the second stage.
+The second stage further refines facial details, producing results that more closely resemble the reference image.
+Zero image guidance.
+Zero image guidance is introduced specifically to handle the occasional color desaturation observed in the edited videos.
+By adjusting the negative prompt to steer the generation away from the zero image direction, this mechanism effectively enhances color fidelity.
+Figure 13 provides an extension of the results in Figure 9 by presenting edits across a wider range of $\gamma$ values.
+As $\gamma$ approaches 1, the contribution of zero image guidance within the negative image guidance increases, leading to noticeably improved color saturation in the edited videos.
+Figure 12: Full qualitative results of the ablation study on the $\rho$-start sampling.
+We present the complete results corresponding to Fig.
+8 in the main manuscript by adding final edited results.
+As the value of $\rho$ increases from 10 to 40 in Stage 1, the parrot in the source video gradually transforms to more closely resemble the owl in the reference image.
+After processing through Stage 2, the final edits further align with the reference appearance and exhibit improved visual quality.
+Figure 13: Full qualitative results of the ablation study on the zero image guidance scale $\gamma$ .
+We include the complete results corresponding to Fig.
+9 in the main paper.
+As the $\gamma$ value is adjusted, the color fidelity of the edited video is calibrated to better match the reference image.
+Figure 14: Example of failure case.
+A large semantic gap between the object in the source video and the reference image poses challenges for our method.
+We show results from two versions of our method, obtained by varying the sampling start point $\rho$ and the dilation kernel size.
+In this case, editing a bird into a red panda results in missing or incorrect transfer of the wings and arms.
+Appendix C Related work
+C.1 Image and video generation
+Recent advancements in latent diffusion models (LDMs) have significantly improved the fidelity and controllability of image generation.
+Text-to-image (T2I) models such as Stable Diffusion [stablediffusion], DALL-E 2 [dalle2], and Imagen [imagen] achieve high-resolution image synthesis aligned with complex text prompts.
+These models operate in the latent space for computational efficiency and support diverse conditioning mechanisms, including text, layout, and edge maps [ipadapter, controlnet].
+Their flexibility has made LDMs foundational components for extending generative modeling into the video domain.
+Motivated by T2I success, text-to-video (T2V) models have emerged to generate coherent video sequences conditioned on textual descriptions.
+CogVideo [cogvideo] adapts T2I backbones with temporal attention, enabling efficient video generation with minimal overhead.
+ZeroScope [zeroscope], built on ModelScope [modelscope], improves visual fidelity and consistency, producing higher-resolution outputs while mitigating common artifacts in earlier methods.
+AnimateDiff [animatediff] integrates temporal modules into Stable Diffusion [stablediffusion], facilitating the T2V model with improved coherence between frames.
+Alongside architectural advances, controllability and identity preservation have gained importance.
+Several methods leverage a reference frame along with a prompt to guide spatial appearance and layout across time [stiv, ti2v, controlvideo].
+In addition, others introduce subject-driven personalization via instance finetuning [videoalchemist, dreamvideo, tune_a_video, customcrafter], enabling video generation with identity consistency across diverse prompts.
+These trends illustrate a broader shift toward multimodal conditioning across text, image, and motion to allow more expressive and controllable video synthesis.
+C.2 Image and video editing
+Following the advances in image and video generation models, diffusion models have also made significant progress in image and video editing, allowing precise content manipulation.
+In the image domain, Prompt-to-Prompt [prompt-to-prompt] introduces cross-attention manipulation within pretrained diffusion models, enabling semantic edits without finetuning or masks.
+InstructPix2Pix [instructpix2pix] leverages pairs of instructions and images to support editing guided by natural language.
+EditAnything [editanything] generalizes this paradigm by integrating multi-modal inputs such as text, masks, and user clicks.
+Video editing extends these capabilities to the temporal dimension.
+Tune-A-Video [tune_a_video] employs one-shot tuning of pretrained image diffusion models for video editing tasks, allowing users to modify video content and style with minimal data and strong temporal consistency.
+Video-P2P [videop2p] extends cross-attention control techniques from image to video, maintaining both spatial and temporal coherence.
+Text2Video-Zero [text2video] removes the need for paired text-video data, instead leveraging pretrained text-image models for flexible and data-efficient video editing.
+To overcome the limitations of editing based solely on text, recent studies propose image-conditioned video editing methods that use a reference image to guide appearance and identity [genvideo, make_a_pro, cutandpaste].
+Make-A-Protagonist [make_a_pro] introduces a unified framework for protagonist replacement, background editing, and style transfer using mask-guided denoising.
+GenVideo [genvideo] proposes shape-aware masking and latent noise correction to ensure temporal consistency and subject fidelity.
+VideoSwap [videoswap] adds semantic point correspondences to provide interactive and precise control over the subject’s motion and appearance.
+Despite these advancements, existing methods still require finetuning on each source video and rely on user-defined semantic keypoints, limiting practicality and scalability.
+In contrast, our method directly utilizes the image prompt without any finetuning, while effectively preserving temporal consistency and semantic accuracy.
+Appendix D Limitation
+Although our method exhibits notable improvements over existing baselines, as evidenced by extensive experiments, it still faces certain limitations.
+When there is a substantial disparity in the innate physical characteristics between the source and target objects, our method struggles to perform object replacement.
+Figure 14 shows that our method generates wings for the red panda which is an anatomically implausible feature to transfer the parrot’s flying motion.
+Conversely, in attempting to preserve the red panda’s natural characteristics, it fails to convey the intended motion.
+Appendix E Additional experimental settings
+Appendix E.1 provides further implementation details, including hyperparameter settings.
+Appendix E.2 outlines the reproducibility settings for all baseline models used in our experiments.
+Appendix E.3 presents the evaluation questions used in our user study.
+E.1 Additional implementation details
+Our method introduces four key hyperparameters: $\rho$, $\gamma$, dilation kernel size, and guidance scale.
+Each parameter has a well-defined and narrow range that yields consistent performance across diverse editing tasks.
+Consequently, hyperparameter tuning in the proposed method requires minimal manual adjustment and can be efficiently performed without extensive search or optimization.
+  • $\rho\in[10,30]$: Controls the starting timestep for the denoising trajectory.
+In practice, one of the discrete values (10, 20, or 30) generally performs well.
+Users may optionally refine the result by exploring intermediate values.
+  • $\gamma\in[0,1]$: Controls the weighted combination of standard zero-vector and zero-image embeddings in zero image guidance.
+The default value of 0.5 has been empirically found to perform robustly across a wide variety of scenes, object types, and motion dynamics.
+  • dilation kernel size $k$ $\in[0,10]$: Defines the kernel size used in the dilated dual masking.
+It affects the spatial smoothness and coverage of the edited region.
+The default value of 3 works well in most cases, while slightly larger values (e.g., 5 or 7) can help when the source and target objects differ significantly in shape or size.
+  • guidance scale $\in[6,10]$: Determines the strength of classifier-free guidance during denoising. The default setting of 6 typically yields semantically faithful results.
+Depending on the resolution of the source video, we use one of the following sizes: 320×576, 432×768, 480×480, 515×512, 600×600, 640×640, or 700×700.
+As an exception, the results in Fig.
+LABEL:fig:teaser are rendered at 1024×576 to better showcase the our method’s performance.
+The number of frames is set to 8 or 16, selected according to the source video length.
+E.2 Reproducibility of baseline models
+All baseline models are run using the official implementations released in their GitHub repositories.
+For a fair comparison, we evaluate all baselines under the same experimental settings as our method unless specified otherwise (refer to Appendix E.1).
+The same text prompts are applied across all methods, except for the method that follows its own text prompting strategy.
+We also use the same spatial resolution for VINO and all baseline models, except for methods that inherently require square inputs.
+  • SDEdit-Video [sdedit] [^1] https://github.com/huggingface/diffusers/blob/main/src/diffusers/pipelines/animatediff/pipeline_animatediff_video2video.py We adopt the video-extended implementation provided by HuggingFace, as the original SDEdit is originally image-based.
+  • FateZero [fatezero] [^2] https://github.com/ChenyangQiQi/FateZero As FateZero [fatezero] supports only square outputs, we paste its generated video into the corresponding region of the source video for comparison.
+Both FateZero [fatezero] and Make-A-Pro [make_a_pro] employ model-based prompting procedures to automatically generate text prompts.
+While FateZero originally uses an earlier BLIP variant for this step, the overall prompting procedure is fundamentally the same as that of Make-A-Protagonist, which uses BLIP-2.
+To maintain consistency, we adopt the more recent BLIP-2–based prompting strategy for FateZero as well and use identical text prompts for both methods.
+  • Video-P2P [videop2p] [^3] https://github.com/dvlab-research/Video-P2P Similar to FateZero, since Video-P2P [videop2p] produces square outputs, we paste its generated video into the corresponding region of the source video.
+  • TokenFlow [tokenflow] [^4] https://github.com/omerbt/TokenFlow TokenFlow is evaluated under the same experimental settings as our method without any additional modifications.
+To implement TokenFlow+Ours, we extend TokenFlow by adding IP-Adapter [ipadapter] to support image conditioning, and incorporate our proposed approaches, including zero-image guidance, $\rho$-start sampling, and dilated dual masking.
+  • AniamteDiff [animatediff] [^5] https://github.com/guoyww/AnimateDiff+IP-Adapter [ipadapter] [^6] https://github.com/tencent-ailab/IP-Adapter For both AnimateDiff and IP-Adapter, we use their official GitHub implementations, and fuse the two following the integration scheme proposed in the IP-Adapter paper.
+  • Make-A-Protagonist [make_a_pro] [^7] https://github.com/HeliosZhao/Make-A-Protagonist Following the original prompt generation strategy, we extract protagonist cues using BLIP-2 [blip2] and generate prompts via visual question answering as provided in its official GitHub repository.
+E.3 User study details
+In the user study described in Section 3.1, participants were asked to rank the seven generated videos from 1st (best) to 7th (worst) for each of the following five evaluation criteria:
+  • Image Alignment: How well the appearance of the "object in the generated video" matches the "object in the reference image (excluding the background)".
+  • Text Alignment: How well "the content of the generated video" aligns with the "provided text prompt".
     Text prompt: “a polar bear skiing”
-  • Motion Alignment: Rank the videos from 1st to 4th based on how closely the "motion of the object in the generated video" matches the "motion of the object in the source video".
-  • Temporal Consistency: Rank the videos from 1st to 4th based on the "visual consistency of the object and background throughout the edited video".
-  • Overall Quality: Taking all the above aspects into consideration, rank the videos from 1st to 4th based on the overall realism and naturalness.
-Appendix G Algorithm
-The overall video editing procedure of ImEdit is summarized in Algorithm 1.
+  • Motion Alignment: How closely the "motion of the object in the generated video" matches the "motion of the object in the source video".
+  • Temporal Consistency: The degree to which the object and background remain consistent across time.
+  • Overall Quality: The overall realism and naturalness of the edited video, considering all aspects above.
+Appendix F Algorithm
+The overall video editing pipeline of VINO is outlined in Algorithm 1.
 Our training-free method consists of two stages: a rough editing stage that generates an initial object transformation guided by the source mask, and a refinement stage that further improves visual fidelity and temporal coherence using an updated mask derived from the intermediate result.
-Algorithm 1 Training-free video editing procedure of ImEdit
+Algorithm 1 Training-free video editing procedure of VINO
 1:Input: Source video $\mathcal{V}_{s}$, image prompt $\mathcal{R}$, text prompt $\mathcal{T}$, negative text prompt $\bar{\mathcal{T}}$
@@
 3: Prompt feature extraction
-4:${\bm{c}}_{\text{img}}\leftarrow\text{CLIP}_{\text{img}}(\mathcal{R}),{\bm{c}}_{\text{txt}}\leftarrow\text{CLIP}_{\text{txt}}(\mathcal{T})$
-5:$\bar{{\bm{c}}}_{\text{img}}=\gamma\cdot\text{CLIP}_{\text{img}}({\mathbf{I}}_{0})+(1-\gamma)\cdot\mathbf{0}$, $\bar{{\bm{c}}}_{\text{txt}}=\gamma\cdot\text{CLIP}_{\text{txt}}({\mathcal{\bar{T}}}),\text{ where }\gamma\in[0,1]$
+[…]
+[…]
 6: Stage 1: Rough editing
 7: Dilated mask generation
-8:${\bm{M}}_{s}\leftarrow\text{Segment}(\mathcal{V}_{s})$
-9:${\bm{M}}^{\prime}_{s}\leftarrow\text{Dilate}({\bm{M}}_{s})$
+[…]
+[…]
 10: Latent encoding
-11:${\bm{z}}_{0}\leftarrow\mathcal{E}(\mathcal{V}_{s})$
-12:$\{{\bm{z}}_{t}\}_{t=\tau_{1}}^{\tau_{\nu}}\leftarrow\text{DDIM-inv}({\bm{z}}_{0})$
+[…]
+[…]
 13: Denoising
 14:Select $\rho\in[\nu]$
-15:${\bm{z}}^{\prime}_{\tau_{\rho}}\leftarrow\sqrt{\alpha_{\tau_{\rho}}}\cdot{\bm{z}}_{0}+\sqrt{1-\alpha_{\tau_{\rho}}}\cdot\eta$, where $\eta\sim\mathcal{N}(0,{\bm{I}})$
-16:for $t=\rho$ to $1$ do
+[…]
+16:for $t=\rho$ to 11 do
 17:  ${\bm{z}}^{\text{{fg}}}_{\tau_{t\texttt{-}1}}\leftarrow\text{DDIM-samp}({\bm{z}}^{\prime}_{\tau_{t}},{\bm{c}}_{\text{img}},{\bm{c}}_{\text{txt}},\bar{{\bm{c}}}_{\text{img}},\bar{{\bm{c}}}_{\text{txt}},\tau_{t})$
@@
 21: Latent decoding
-22:$\mathcal{V}_{r}\leftarrow\mathcal{D}({\bm{z}}^{\prime}_{\tau_{0}})$
-23:$\mathcal{V}_{r}\leftarrow\mathcal{V}_{r}\odot{\bm{M}}^{\prime}_{s}+\mathcal{V}_{s}\odot(1-{\bm{M}}^{\prime}_{s})$
+[…]
+[…]
 24: Stage 2: Refined editing
 25: Dilated dual mask generation
-26:${\bm{M}}_{r}\leftarrow\text{Segment}(\mathcal{V}_{r})$
-27:${\bm{M}}^{\prime}_{r}\leftarrow\text{Dilate}({\bm{M}}_{r})$
-28:${\bm{M}}_{e}\leftarrow{\bm{M}}^{\prime}_{s}\cup{\bm{M}}^{\prime}_{r}$
+[…]
+[…]
+[…]
 29: Latent encoding
-30:${\bm{z}}_{0}\leftarrow\mathcal{E}(\mathcal{V}_{r})$
-31:$\{{\bm{z}}_{t}\}_{t=\tau_{1}}^{\tau_{\nu}}\leftarrow\text{DDIM-inv}({\bm{z}}_{0})$
+[…]
+[…]
 32: Denoising
 33:Select $\rho\in[\nu]$
-34:${\bm{z}}^{\prime}_{\tau_{\rho}}\leftarrow\sqrt{\alpha_{\tau_{\rho}}}\cdot{\bm{z}}_{0}+\sqrt{1-\alpha_{\tau_{\rho}}}\cdot\eta$, where $\eta\sim\mathcal{N}(0,{\bm{I}})$
-35:for $t=\rho$ to $1$ do
+[…]
+35:for $t=\rho$ to 11 do
 36:  ${\bm{z}}^{\text{{fg}}}_{\tau_{t\texttt{-}1}}\leftarrow\text{DDIM-samp}({\bm{z}}^{\prime}_{\tau_{t}},{\bm{c}}_{\text{img}},{\bm{c}}_{\text{txt}},\bar{{\bm{c}}}_{\text{img}},\bar{{\bm{c}}}_{\text{txt}},\tau_{t})$
@@
 40: Latent decoding
-41:$\mathcal{V}_{e}\leftarrow\mathcal{D}({\bm{z}}^{\prime}_{\tau_{0}})$
-Appendix H Additional qualitative results
+[…]
+Appendix G Additional qualitative results
+Additional qualitative comparisons with baseline methods are provided in Fig.
+15, Fig.
+16, Fig.
+17, Fig.
+18, and Fig.
+19, which further demonstrate the superior visual fidelity and temporal consistency achieved by our approach.
 Notably, AnimateDiff+IP-Adapter frequently exhibits noticeable distortions in both shape and color, leading to unnatural visual artifacts.
@@
 Figure 15: Complete video editing comparisons with baseline methods.
-FateZero fatezero and Make-a-Pro make_a_pro use text prompts generated by BLIP-2 blip2 , following their respective proposed strategies.
-Video-P2P videop2p and FateZero fatezero use only text prompts, while Make-a-Pro make_a_pro , AnimateDiff animatediff +IP-Adapter ipadapter , and our proposed method use both image and text prompts.
+FateZero [fatezero] and Make-a-Pro [make_a_pro] use text prompts generated by BLIP-2 [blip2], following their respective proposed strategies.
+TokenFlow [tokenflow], Video-P2P [videop2p], FateZero [fatezero], and SDEdit-Video [sdedit] use only text prompts, while Make-a-Pro [make_a_pro], AnimateDiff [animatediff]+IP-Adapter [ipadapter], and our proposed method use both image and text prompts.
 Figure 16: Complete video editing comparisons with baseline methods.
-FateZero fatezero and Make-a-Pro make_a_pro use text prompts generated by BLIP-2 blip2 , following their respective proposed strategies.
-Video-P2P videop2p and FateZero fatezero use only text prompts, while Make-a-Pro make_a_pro , AnimateDiff animatediff +IP-Adapter ipadapter , and our proposed method use both image and text prompts.
+FateZero [fatezero] and Make-a-Pro [make_a_pro] use text prompts generated by BLIP-2 [blip2] following their respective proposed strategies.
+TokenFlow [tokenflow], Video-P2P [videop2p], FateZero [fatezero], and SDEdit-Video [sdedit] use only text prompts, while Make-a-Pro [make_a_pro], AnimateDiff [animatediff]+IP-Adapter [ipadapter], and our proposed method use both image and text prompts.
 Figure 17: Complete video editing comparisons with baseline methods.
-FateZero fatezero and Make-a-Pro make_a_pro use text prompts generated by BLIP-2 blip2 following their respective proposed strategies.
-Video-P2P videop2p and FateZero fatezero use only text prompts, while Make-a-Pro make_a_pro , AnimateDiff animatediff +IP-Adapter ipadapter , and our proposed method use both image and text prompts.
+FateZero [fatezero] and Make-a-Pro [make_a_pro] use text prompts generated by BLIP-2 [blip2] following their respective proposed strategies.
+TokenFlow [tokenflow], Video-P2P [videop2p], FateZero [fatezero], and SDEdit-Video [sdedit] use only text prompts, while Make-a-Pro [make_a_pro], AnimateDiff [animatediff]+IP-Adapter [ipadapter], and our proposed method use both image and text prompts.
+Figure 18: Complete video editing comparisons with baseline methods.
+FateZero [fatezero] and Make-a-Pro [make_a_pro] use text prompts generated by BLIP-2 [blip2] following their respective proposed strategies.
+TokenFlow [tokenflow], Video-P2P [videop2p], FateZero [fatezero], and SDEdit-Video [sdedit] use only text prompts, while Make-a-Pro [make_a_pro], AnimateDiff [animatediff]+IP-Adapter [ipadapter], and our proposed method use both image and text prompts.
+Figure 19: Complete video editing comparisons with baseline methods.
+FateZero [fatezero] and Make-a-Pro [make_a_pro] use text prompts generated by BLIP-2 [blip2] following their respective proposed strategies.
+TokenFlow [tokenflow], Video-P2P [videop2p], FateZero [fatezero], and SDEdit-Video [sdedit] use only text prompts, while Make-a-Pro [make_a_pro], AnimateDiff [animatediff]+IP-Adapter [ipadapter], and our proposed method use both image and text prompts.